### v1

[…]
Advances in computational power and AI have increased interest in reinforcement learning approaches to inventory management. This paper provides a theoretical foundation for these approaches and investigates the benefits of restricting to policy structures that are well-established by decades of inventory theory. In particular, we prove generalization guarantees for learning several well-known classes of inventory policies, including base-stock and $(s,S)$ policies, by leveraging the celebrated Vapnik-Chervonenkis (VC) theory. We apply the concepts of the Pseudo-dimension and Fat-shattering dimension from VC theory to determine the generalizability of inventory policies, that is, the difference between an inventory policy’s performance on training data and its expected performance on unseen data. We focus on a classical setting without contexts, but allow for an arbitrary distribution over demand sequences and do not make any assumptions such as independence over time. We corroborate our supervised learning results using numerical simulations.
Managerially, our theory and simulations translate to the following insights. First, there is a principle of ?learning less is more? in inventory management: depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies to minimize overfitting errors. Second, the number of parameters in a policy class may not be the correct measure of overfitting error: in fact, the class of policies defined by $T$ time-varying base-stock levels exhibits a generalization error comparable to that of the two-parameter $(s,S)$ policy class. Finally, our research suggests situations in which it could be beneficial to incorporate the concepts of base-stock and inventory position into black-box learning machines, instead of having these machines directly learn the order quantity actions.
[…]
VC theory, Pseudo-dimension, data-driven algorithm design, generalization bound, estimation error, sample complexity, shattering, inventory, base-stock policy, ($s,S$) policy, less is more
[…]
May 1, 2024
But it is the VC dimension rather than the number of parameters of the set of functions that defines the generalization ability of a learning machine.(Vapnik 1998, Page 160)
1 Introduction
Inventory management plays a crucial role in businesses, ranging from manufacturing to distribution to retailing, and is arguably the most fundamental topic in operations and supply chain management. The design of effective inventory algorithms is particularly important in the modern era, characterized by the rapid proliferation of AI technology, because an increasingly significant portion of inventory replenishment decisions has shifted from human intuition to algorithmic decision-making (e.g., Liu et al. 2023). A classical approach to inventory management involves estimating a stochastic process for demand over time based on past data and optimizing around this estimate. However, model selection is notoriously difficult, and it is also unclear whether the two-step process of estimating and then optimizing leads to the best decisions (e.g., Elmachtoub and Grigas 2022).
An emerging alternative approach to inventory management involves learning policies directly from past data. Indeed, under the assumption that demands are uncensored—a common assumption in inventory literature—it becomes possible to accurately evaluate the counterfactual scenario of how any policy would have performed given any historical demand sequence. This allows for the adoption of supervised learning or a ?hindsight? reinforcement learning approach (Sinclair et al. 2023), which has been demonstrated to be computationally feasible in terms of policy differentiation (Alvo et al. 2023) and has even been deployed in practice (Madeka et al. 2022). The latter papers highlight the power of black-box neural networks, which do not require explicit knowledge about the structure of optimal inventory policies to achieve high performance.
By contrast, our paper demonstrates the potential benefits of adhering to policy structures that have been well-established in 70 years of inventory theory, beginning with Arrow et al. (1951). We derive learning theory results and complement them with simulations. We analyze the fundamental classes of base-stock and $(s,S)$ policies, which are optimal in various settings with independent demand over time (e.g., Snyder and Shen 2019). We approach this as a supervised learning problem over demand sequences, in the spirit of Madeka et al. (2022), Alvo et al. (2023), but focus on a classical setting without contexts. We use the celebrated Vapnik-Chervonenkis (VC) theory to bound the learning-theoretic complexity of these policy classes when trained offline on past demand sequences, without assuming they are independent over time as this is often impossible to verify in practice. We provide statistical learning guarantees for this training procedure, which resembles the ?hindsight? reinforcement learning approaches mentioned earlier.
VC theory is a foundational framework in the field of statistical learning theory (Vapnik and Chervonenkis 1971). At a high level, VC theory helps us understand the complexity of a hypothesis class in terms of its ability to generalize to unseen data, which is critical for avoiding overfitting. The core concept of VC theory centers on the VC dimension, along with its generalizations and variations, such as the Pseudo-dimension (Pollard 1984), and the Fat-shattering dimension, also known as the Pseudoγ-dimension (Kearns and Schapire 1994). These dimensions can be used to provide a distribution-free upper bound on the Rademacher complexity, and overall, quantify the capability of a hypothesis class to learn complex patterns. A hypothesis class with a higher dimension generally possesses a greater potential to learn a more complex pattern from data, but may suffer from reduced generalizability due to overfitting, leading to the need for a larger training dataset to ensure good performance of the learned hypothesis.
In the context of inventory management, VC theory can be a powerful tool for establishing rigorous bounds on the generalization error of inventory policy classes, namely, the difference between an inventory policy’s performance on training data and its expected performance on true data. This is useful because, as we will show, naive metrics such as the number of parameters do not necessarily indicate generalizability. We instead use VC theory to indicate generalizability, and corroborate our theoretical results using simulations. All in all, this reveals how effective each of these policy class restrictions is in terms of reducing overfitting, complementing the large body of inventory theory that already characterizes when each of these policy classes are optimal.
1.1 ?Learning Less is More? Phenomenon
Let us first illustrate why restricting the inventory policies to a suboptimal class can be desirable. Consider a classical backlogged inventory model with fixed costs and zero lead time. It is well known that the optimal policy has an ($s,S$) structure: when the inventory level reaches the reorder point $s$, an order is placed to bring the inventory level up to the order-up-to level $S$. Given substantial historical data on demand, it might be plausible to learn the optimal ($s^{*},S^{*}$) parameters for the true demand distribution. However, in scenarios where historical data is scant, learning the optimal $(s,S)$ parameters for the observed demands can lead to overfitting. In such scenarios, restricting to simpler policy classes like base-stock policies, where $s=S$, might exhibit superior performance compared to the more intricate ($s,S$) policies. We illustrate this idea conceptually in Figure 1, which is supported by our theoretical and simulation results in the paper. Figure 1 provides important managerial insights into the principle of ?learning less is more? in inventory management: depending on the amount of data available, we may want to restrict the policy space.
Figure 1: This figure conceptually demonstrates the ?learning less is more? phenomenon in the context of inventory management. It suggests that in scenarios with limited historical data, learning a simpler base-stock policy is more advantageous than attempting to learn the more complex optimal ($s,S$) policy, even though the latter is required for optimality due to fixed costs. As the volume of available data increases, it becomes appropriate to transition towards learning the more complex ($s,S$) policy.
The phrase ?less is more,? as coined by Theodore R. Sizer in his book ?Horace’s Compromise? (Sizer 1984), embodies one of the Coalition’s Nine Common Principles aimed at guiding school reform initiatives. According to Sizer, the aphorism highlights the importance of students engaging deeply and mastering a few subjects, rather than superficially skimming through a broader range of topics. The principle of less is more has also been investigated and analyzed in machine learning, sometimes using structural risk minimization (e.g., Chapter 7.2 of Shalev-Shwartz and Ben-David 2014), which balances between in-sample performance and model simplicity. In the context of inventory management, the principle of ?learning less is more? suggests that the employment of more sophisticated algorithms may lead to analysis paralysis, and the additional algorithmic complexity does not necessarily translate to better decisions.
Mirroring the educational process undertaken with students, we aim to teach algorithms effective inventory decision-making. We adopt the ?learning less is more? philosophy and provide generalization guarantees for simple inventory policies trained on demand sequences by applying VC theory. Interestingly, we theoretically prove that the generalization error of ($s,S$) policies is an order of magnitude higher than that of base-stock policies. This explains why learning a base-stock policy with one parameter can be preferable to learning an ($s,S$) policy with two parameters. Overall, our results provide strong theoretical support for the phenomenon observed in Figure 1.
On the other hand, our results also reveal a surprisingly mild manifestation of the ?learning less is more? phenomenon in other settings. Consider $(S^{t})$ policies, another generalization of the single-parameter base-stock policies, where we now allow varying base-stock levels $\{S^{t}\}_{t=1}^{T}$ to accommodate non-stationary demand over a $T$-period. We show that the generalization error of $(S^{t})$ policies is in fact of the same order as base-stock policies, despite having $T$ parameters. Numerically, we observe that the generalization error of $(S^{t})$ policies is comparable to that of $(s,S)$ policies, despite having significantly more parameters.
1.2 Outline of Results
We consider inventory replenishment for a single durable good over a finite, discrete time horizon of length $T$. We assume backlogged demand and allow for a deterministic lead time. We define performance by the average inventory cost over the time horizon, which includes fixed costs for each replenishment, backlogging costs, and holding costs. Demands, cost parameters, and policy parameters are normalized so that given any $T$, the performance of any policy on any demand realization is bounded by a constant. The demand realization indicates the demand for the good during each time period, drawn from an arbitrary distribution over sequences of length $T$ (that could exhibit correlations over time). We assume access to $N$ independent and identically distributed (i.i.d.) samples of sequences drawn from the same distribution, and investigate how the uniform generalization error for a policy class $\Pi$ (difference between expected performance on the true distribution, and average empirical performance over the $N$ samples, taking the supremum over policies in $\Pi$) scales with $N$ and $T$. We consider the three different policy classes for $\Pi$: base-stock policies, $(s,S)$ policies, and $(S^{t})$ policies.
We derive the following results:
  1. 1.
    Base-stock policies (or ?$S$ policies?): expected generalization error is $\Theta(\sqrt{1/N})$;
  2. 2.
    $(s,S)$ policies: expected generalization error is $O(\sqrt{\log T/N})$ and $\Omega(\sqrt{(\log T/\log\log T)/N})$;
  3. 3.
    $(S^{t})$ policies: expected generalization error is $\Theta(\sqrt{1/N})$.
We note that our result for $(S^{t})$ policies assumes no fixed replenishment costs, which is standard whenever $(S^{t})$ policies are analyzed. Under this assumption, $(S^{t})$ policies have a generalization error that is constant in $T$, while having $T$ parameters. In stark contrast, $(s,S)$ policies have a generalization error that is growing with $T$, despite having only two parameters (we also note that our lower bound for $(s,S)$ policies holds even when there are no fixed replenishment costs). Costs aside, this provides fundamental intuition about the volatility of state trajectories of different policy classes, suggesting that the number of parameters is not the right measure of complexity. Indeed, perturbing the parameters $(s,S)$ can cause a missed inventory replenishment and substantial divergence in the future state trajectory that persists over a long $T$. In particular, a small perturbation to the parameter $s$ or $S$ can completely change the set of time periods during which inventory is replenished. By contrast, the state trajectories of $(S^{t})$ policies are ?stable? in the right measure—exhibiting a Pseudoγ-dimension that does not grow with $T$. Intuitively, a small perturbation to any parameter $S^{t}$ can only slightly change the inventory level.
Our results can also be stated in equivalent alternative forms, such as sample complexity. In this form, our results say that to guarantee an expected generalization error of at most $\epsilon$, the number of samples required is growing with the horizon length $T$ for $(s,S)$ policies, but not growing with $T$ for base-stock or $(S^{t})$ policies. All of these expectation bounds can also be converted into high probability bounds using standard concentration inequalities.
Our theoretical results require a problem-specific analysis of inventory trajectories over time, as needed to bound Pseudo-dimension and Pseudoγ-dimension, which to our knowledge is new to both inventory theory and learning theory. We also emphasize that even if one assumes each period to have its own independent demand distribution, results 2. and 3. above were not previous known. In particular, existing works (e.g., Levi et al. 2007, Cheung and Simchi-Levi 2019) have analyzed the sample complexity of $(S^{t})$ policies under the independence assumption (in which case they are optimal) and shown it to grow with $T$, in a related ?multiplicative? setting. Therefore, we believe the non-dependence on $T$ in our additive setting to be quite surprising. We elaborate on these points in Section 4.1.
Finally, we show that the aforementioned surprising distinction between $(s,S)$ and $(S^{t})$ policies suggested by our worst-case theoretical guarantees do translate to milder numerical settings. We consider $(s,S)$ policies in a setting with fixed cost (where they are near-optimal), and $(S^{t})$ policies in a setting without fixed cost but with non-stationary independent demand (where they are optimal). In both settings we compare the out-of-sample error to that of ?learning less?: learning just a base-stock policy. Overall, the intersection point in Figure 1 emerges further to the right for $(s,S)$ policies than $(S^{t})$ policies, supporting our theoretical finding that a manager needs more data samples before they should want to ?learn more? in the $(s,S)$ setting.
1.3 The Value of Inventory Theory
Section 1.1 discussed how restricting to suboptimal inventory policy classes can reduce overfitting, and Section 1.2 explained how the number of parameters is not the right indicator of overfitting. Finally, we would like to distinguish between our approach of learning policy parameters (e.g., $s,S,S^{t}$), and learning order quantities (i.e., ?actions?) which is the prevailing approach in papers that use state-of-the-art deep reinforcement learning for inventory management (e.g., Gijsbrechts et al. 2022, Madeka et al. 2022, Oroojlooyjadid et al. 2022, Alvo et al. 2023, Liu et al. 2023, Qi et al. 2023). Although these papers focus on engineering realistic contextual settings, whereas we focus on a parsimonious theoretical setting, the comparison between learning policy parameters and learning order quantities can be made irrespective of the presence of contexts.
We make this comparison numerically in a concise non-contextual setting in Section 5.3. We consider zero fixed replenishment costs, a short positive lead time, and i.i.d. demands. We first assume that demands are backlogged, and compare learning a base-stock policy, which is provably optimal from inventory theory, to learning an arbitrary function that maps inventory state to order quantity and is agnostic to inventory theory. Our experiments show that learning a base-stock policy is much more sample-efficient. Next, we assume that demands are lost, in which case a base-stock policy is no longer optimal, whereas the arbitrary function can capture the optimal policy. Nonetheless, our results still demonstrate a significant advantage in learning a base-stock policy at small sample sizes, i.e., ?learning less is more.? We emphasize that in these experiments, the lead time is very short, hence the state space to be captured by the arbitrary function is small, which should only understate the benefits of learning less.
Although not analyzed in this paper, our approach of learning policy parameters can be engineered in a contextual setting. To elaborate, at a time $t$, let $\textbf{X}_{t}$ denote the context vector including all information about upcoming demand, and $\textbf{I}_{t}$ denote the vector of on-hand/pipeline inventory. The aforementioned black-box learning approaches would learn an order quantity action $a(\textbf{X}_{t},\textbf{I}_{t})$ taking both $\textbf{X}_{t}$ and $\textbf{I}_{t}$ into the state, whereas we would learn a desired base-stock level $S(\textbf{X}_{t})$ and force the order quantity $a(\textbf{X}_{t},\textbf{I}_{t})$ to be $S(\textbf{X}_{t})$ minus the total inventory in $\textbf{I}_{t}$. Of course, we are merely restricting ourselves to a special functional form of their approach, regarding how $\textbf{I}_{t}$ should impact $a(\textbf{X}_{t},\textbf{I}_{t})$, which can sometimes be rediscovered by the black-box approach given enough data (Alvo et al. 2023). Nonetheless, it would be interesting to further investigate how this restriction could potentially improve performance, especially when relevant data is limited. Our hope is that by combining inventory concepts such as ?base-stock? $S$ and ?reorder point? $s$ with machine learning algorithms to handle high-dimensional contexts $\textbf{X}_{t}$, classical inventory theory and modern artificial intelligence can co-exist in harmony.
1.4 Organization of Paper
The rest of the paper is organized as follows. We start by discussing the related work in Section 2. We describe our model setup in Section 3, and define estimation error and approximation error in Sections 3.1 and 3.2, respectively. We review several well-known results from learning theory by using the concepts of Rademacher complexity, VC-dimension, Pseudo-dimension, and Pseudoγ-dimension in Section 3.3. We present our theoretical results on distribution-free bounds for the expected generalization errors of three parametric policy classes in Section 4. Numerical experiments are presented in Section 5 and we conclude the paper in Section 6.
2 Further Related Work
Data-driven inventory management: Most papers on data-driven inventory management focus on the Newsvendor model (e.g., Levi et al. 2015, Lin et al. 2022, Besbes and Mouchtaki 2023, Besbes et al. 2023), with some exceptions. Levi et al. (2007) and Cheung and Simchi-Levi (2019) both consider dynamic models and prove sample complexity bounds on base-stock policies under the assumption that demands are independent across time but not necessarily identically distributed. Huh and Rusmevichientong (2009) use stochastic gradient descent to learn inventory policies for demand that is both independent and identically distributed over time. Ban (2020) considers a dynamic model with demand censoring and proves the consistency of a non-parametric estimator of $(s_{t},S_{t})$. Yuan et al. (2021) provide an upper bound on the regret of learning the optimal $(s,S)$ policy.
It is quite surprising that the study of VC-type complexity, even for the simplest classes such as base-stock and ($s,S$) policies, is missing; this is the gap we aim to fill with this paper. It is worth noting that Han et al. (2023) apply existing bounds on the Pseudo-dimension of neural networks directly to solve a Newsvendor problem. Their model is based on the contextual Newsvendor model that was pioneered by Ban and Rudin (2019). By contrast, our analysis focuses on the dynamics of inventory state transitions over time and how this relates to learning-theoretic dimensions. We note that recently, Guan and Mišić (2022) have asked similar questions for optimal stopping, which also considers dynamic decision-making over time but is otherwise unrelated to inventory.
The ?learning less is more? phenomenon advocates for simplicity. In inventory management, this phenomenon has been observed in Lyu et al. (2024), when designing UCB-type algorithms in a lost-sales model with lead times, to learn the optimal base-stock policy with one parameter and the optimal capped base-stock policy with two parameters. They note that learning the optimal one-parameter policy is relatively easier, whereas learning the optimal two-parameter policy is more challenging, despite being potentially more rewarding. However, their finds are purely numerical. To the best of our knowledge, our paper is the first one developing a theoretic framework to formally analyze this phenomenon in the context of data-driven inventory management.
Data-driven algorithm design: We are tuning the parameters of policy classes based on historical sequences (instances), a process known as data-driven algorithm design (Balcan 2020, Gupta and Roughgarden 2020). However, we derive problem-specific results by analyzing the sensitivity of inventory state transitions to policy parameters, which appear difficult to deduce from general theory (e.g., Balcan et al. 2021). We also believe that the inventory problem, with its combination of features—dynamic decision-making, state transitions, and distributions over arbitrary sequences—is rather underexplored in this area, which could inspire future data-driven algorithm design for other online algorithms problems. We note that online algorithms can be tuned based on past arrival sequences by using the so-called ?Algorithms with Predictions? framework (Lykouris and Vassilvitskii 2021); however, the theoretical results therein are of a completely different nature.
3 Model and Preliminaries
In this paper, we consider a discrete-time backlogged inventory model with lead time over a finite-horizon. There is a lead time $L\in\mathbb{Z}_{\geq 0}$ between when an order is placed and when it is delivered. There are $T+L$ periods in the entire horizon. The demand process follows a stochastic process with distribution $\mathcal{D}$, which can be non-stationary or correlated across time. A realized demand sequence $\mathbf{d}\triangleq(d^{1},\ldots,d^{T+L})$ consists of demand $d^{t}$ for period $t=1,\ldots,T+L$. Let $h$ and $b$ be the per-unit holding and backlogging costs, respectively, and let $K$ be the fixed ordering cost. Without loss of generality, we normalize $h,b,K\in[0,1]$ and assume $\mathcal{D}$ is supported on $[0,U]^{T+L}$ for a given constant $U$. The decision-maker must specify an inventory control policy $\pi$ that determines an order quantity at each time point with the objective of minimizing the expected average loss over the entire horizon.
Let $x^{t}$ be the inventory level at the beginning of period $t$. Let $(q^{t-L},\ldots,q^{t-1})$ be the pipeline inventory at the beginning of period $t$, where $q^{t}$ denotes the inventory ordered in period $t$ that is to arrive in period $t+L$. We assume empty initial pipeline: $q^{t}=0$ for $t\leq 0$. Let $I^{t}$ be the inventory position at the beginning of period $t$ before ordering, namely, $I^{t}=x^{t}+\sum\limits_{t^{\prime}=t-L}^{t-1}q^{t^{\prime}}$. Let $y^{t}$ be the inventory level after replenishment but before seeing demand in period $t$, namely, $y^{t}=x^{t}+q^{t-L}$. Demands are backlogged and the inventory level in period $t+1$ is updated according to $x^{t+1}=y^{t}-d^{t}$. The sequence of events in period $t$ is summarized as follows:
  1. 1.
    The decision-maker observes the inventory level $x^{t}$ and pipeline vector $\left(q^{t-L},\ldots,q^{t-1}\right)$;
  2. 2.
    A new order $q^{t}$ is placed based on policy $\pi$;
  3. 3.
    The inventory level is updated according to $y^{t}=x^{t}+q^{t-L}$;
  4. 4.
    Demand $d^{t}$ is realized and the loss $\ell^{t}(\pi,\mathbf{d})$ is incurred.
In particular, define $c(x)\triangleq h[x]^{+}+b[-x]^{+}$, where $[\cdot]^{+}\triangleq\max\{\cdot,0\}$, and the loss $\ell^{t}(\pi,\mathbf{d})\triangleq c(y^{t}-d^{t})+K\cdot{\mathbb{I}}\{q^{t-L}>0\}$. Here, $\ell^{t}$ includes the holding and backlogging costs in period $t$, as well as the fixed ordering cost associated with the order placed in period $t-L$ that is arriving during period $t$. Define $\ell(\pi,\mathbf{d})\triangleq\frac{1}{T}\sum\limits_{t=L+1}^{T+L}\ell^{t}(\pi,\mathbf{d})$ as the losses from period $L+1$ to period $T+L$ of a policy $\pi$ (averaged over time) on a given demand sequence $\mathbf{d}$. Here, we ignore the losses associated with periods 1 to $L$ because our first action $q^{1}$ only impacts losses starting from period $L+1$.
For the data-driven problem, the training dataset contains $N$ demand samples $\mathbf{d}_{1},\ldots,\mathbf{d}_{N}$ drawn i.i.d. from $\mathcal{D}$, and each of the samples $\mathbf{d}_{i}\triangleq(d_{i}^{t})_{t=1}^{T+L}$ consists of demand realizations from period $1$ to period $T+L$. Note that this assumption is also made in Ban (2020) and Madeka et al. (2022, §7). Here, the $N$ samples can be interpreted as the demands of $N$ similar products over $(T+L)$ periods, or as $N$ records of one product’s demands over $(T+L)$ consecutive periods (e.g., each sample $i=1,\ldots,N$ represents one year and consists of $T+L$ periods).
3.1 Estimation Error
To take advantage of existing results on inventory management, we aim to learn a policy $\pi$ from a well-studied parametric class of policies denoted by $\Pi$, based on the training dataset $(\mathbf{d}_{i})_{i=1}^{N}$. Note that the family $\Pi$ (to be specified in Section 3.2 below) does not necessarily contain the optimal policy. We define $R(\pi)\triangleq\mathbb{E}_{\mathbf{d}\sim\mathcal{D}}[\ell(\pi,\mathbf{d})]$ as the true risk (i.e., expected loss) of policy $\pi$, and $\hat{R}(\pi)\triangleq\frac{1}{N}\sum_{i=1}^{N}\ell(\pi,\mathbf{d}_{i})$ as its associated empirical risk. Let ${\cal L}(\Pi)\triangleq\{\ell(\pi,\cdot):\pi\in\Pi\}$ denote a set of loss functions induced by the policy class $\Pi$, where each function corresponds to a policy $\pi\in\Pi$. Let $\pi^{*}\in\arg\inf_{\pi\in\Pi}R(\pi)$ denote an optimal policy in the policy class $\Pi$. Given the dataset $(\mathbf{d}_{i})_{i=1}^{N}$, the decision-maker uses a procedure to specify a policy $\pi\in\Pi$. One popular procedure is empirical risk minimization (ERM), which produces an output $\hat{\pi}\in\arg\inf_{\pi\in\Pi}\hat{R}(\pi)$. The question of interest is to bound the expected risk gap between the policy $\hat{\pi}$ produced by ERM and the optimal one $\pi^{*}$ in the policy class $\Pi$. This gap is known as the estimation error (EE) of ERM with respect to (w.r.t.) the function class ${\cal L}(\Pi)$ in the literature (e.g., Section 5.2 of Shalev-Shwartz and Ben-David 2014), denoted by $\mathrm{EE}({\cal L}(\Pi))\triangleq R(\hat{\pi})-R(\pi^{*}).$ The expected estimation error w.r.t. the $N$ samples $(\mathbf{d}_{i})_{i=1}^{N}$ can be bounded from above as follows:
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}[\mathrm{EE}({\cal L}(\Pi))]$ $=\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[R(\hat{\pi})-\hat{R}(\hat{\pi})+\hat{R}(\hat{\pi})-\hat{R}(\pi^{*})+\hat{R}(\pi^{*})-R(\pi^{*})\right]$
$\leq\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\sup_{\pi\in\Pi}\left\{R({\pi})-\hat{R}({\pi})\right\}\right],$ (3.1)
where the inequality comes from the optimality of $\hat{\pi}$ and the fact that $\mathbb{E}\left[\hat{R}(\pi^{*})-R(\pi^{*})\right]=0$. Note that from an upper bound of the expected estimation error, one can directly derive the sample complexity, i.e., the number of samples required to guarantee a given estimation error with a given high probability. The supremum term in (3.1) is referred to as the (uniform) generalization error (GE) or the representativeness of the policy class $\Pi$ in the literature, which describes the worst-case scenario of overfitting within the policy class. We now formally define the generalization error of the policy class $\Pi$ w.r.t. the function class ${\cal L}(\Pi)$:
$\mathrm{GE}({\cal L}(\Pi))\triangleq\sup_{\pi\in\Pi}\left\{R({\pi})-\hat{R}({\pi})\right\}.$ (3.2)
It quantifies the generalization ability of an algorithm (or a policy), namely, the gap between the algorithm’s in-sample performance on the training set and its out-of-sample performance. Note that another widely accepted definition of uniform generalization error is $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$, which is commonly used for deriving high-probability sample complexity bounds. The analysis based on this definition is analogous to that by (3.2), expect for a constant factor.
In this paper, our primary emphasis is on analyzing the expected generalization error w.r.t. the samples $(\mathbf{d}_{i})_{i=1}^{N}$, with the objective of quantifying the learning complexity intrinsic to three different policy classes. More specifically, we aim to upper-bound the generalization error w.r.t. the sample size $N$, which would then lead to an upper bound on the estimation error. Intuitively, with a larger horizon length $T$, the inventory dynamics simulated by the $N$ samples become more unstable and could lead to greater overfitting. We will investigate how the growth rate of the generalization error w.r.t. $T$ varies among different policy classes, by leveraging the concepts of VC-Dimension, Pseudo-Dimension and Pseudoγ-Dimension.
3.2 Approximation Error
The approximation error of a policy class $\Pi$ is defined as the gap between the best policy within the class, $\pi^{*}\in\arg\min_{\pi\in\Pi}R(\pi)$, and the overall optimal policy, $\pi^{\text{opt}}\in\arg\min_{\pi}R(\pi)$. More formally, $\mathrm{AE}({\cal L}(\Pi))\triangleq R(\pi^{*})-R(\pi^{\text{opt}})$. It depends only on the richness of the policy class $\Pi$ and is independent of the sample size $N$.
The gap between the policy $\hat{\pi}$ as determined by the ERM procedure and the overall optimal policy $\pi^{\text{opt}}$ is equal to the sum of the estimation error $\mathrm{EE}({\cal L}(\Pi))$ and the approximation error $\mathrm{AE}({\cal L}(\Pi))$. There is a trade-off in balancing the estimation error and approximation error: a richer policy class results in a lower approximation error but might lead to overfitting and a larger estimation error. How to design policy classes with a balanced trade-off between estimation error and approximation error has been explored in the learning theory literature. The approach involves using prior or domain knowledge about specific problems to initially restrict $\Pi$ with certain structures. Moreover, because the approximation error does not depend on the size of the training data $N$, this independence motivates an analysis of the estimation error w.r.t. $N$ for specific policy classes.
In this paper, we mainly focus on three policy classes:
  • $\Pi_{S}\triangleq\{S:S\in[0,H]\}$: the class of all base-stock policies with a stationary base-stock level $S$. These policies order to bring the inventory position up to $S$ whenever the inventory position falls below $S$. See Section 4.2 for a formal specification.
  • $\Pi_{(s,S)}\triangleq\{(s,S):\underline{H}\leq s\leq S\leq H,S\geq 0\}$: the class of all $(s,S)$ policies. These policies order to bring the inventory position up to $S$ whenever the inventory position is below the reorder point $s$. See Section 4.3 for a formal specification.
  • $\Pi_{(S^{t})}\triangleq\{(S^{t})_{t=1}^{T+L}:(S^{t})_{t=1}^{T+L}\in[0,H]^{T+L}\}$: the class of all base-stock policies with non-stationary levels $(S^{t})_{t=1}^{T+L}$. These policies order to bring the inventory position up to $S^{t}$ when the inventory position falls below $S^{t}$ in period $t$. See Section 4.4 for a formal specification.
All values of $S$ and $S^{t}$ fall within the interval $[0,H]$ for a fixed constant $H>0$, and values of $s$ are lower bounded by $\underline{H}$. We assume $x^{1}=-\infty$ to ensure that a replenishment is always ordered in period $t=1$. In addition, we assume that the parameter domains are continuous, allowing $s,S,S^{t}$ to assume fractional values.
These three classes of policies have been extensively studied in the inventory literature and are widely implemented in practice. We next review several classical results for finite-horizon backlogged inventory models in the stochastic inventory control literature (e.g., Snyder and Shen 2019), demonstrating the conditions under which these policies achieve optimality (i.e., zero approximation error).
  • In settings where the fixed cost $K=0$ and demands are i.i.d. across time, an optimal base-stock policy exists with a stationary base-stock level for each period, falling within the class $\Pi_{S}$ with $H=(L+1)U$.
  • In settings where $K>0$ and demands are i.i.d. across time, an $(s,S)$ policy with both stationary order-up-to level $S$ and reorder point $s$ is asymptotically optimal (as $T\to\infty$), falling within the class $\Pi_{(s,S)}$ where $H$ and $\underline{H}$ are selected appropriately based on the cost parameters (to be further discussed in Section 4.3).
  • In settings where the fixed cost $K=0$ and demands are independently but not necessarily identically distributed across time, an optimal base-stock policy exists. However, this policy features a time-dependent base-stock level for period $t$, which falls within the class $\Pi_{(S^{t})}$ with $H=(L+1)U$.
3.3 Rademacher Complexity, VC-Dimension, Pseudo-Dimension, and Pseudoγ-Dimension
In this section, we review several well-known results from learning theory by using the notion of Rademacher complexity, VC-dimension, Pseudo-dimension, and Pseudoγ-dimension. We will change our notation to generic functions $f(\pi,\cdot)$, which are usually identical to the loss functions $\ell(\pi,\cdot)$. This change provides flexibility, allowing us to use $f(\pi,\cdot)$ to also represent the final inventory levels induced by policies $\pi\in\Pi$. Let $[m]\triangleq\{1,\ldots,m\}$. We first introduce a fundamental concept, known as Rademacher complexity, to bound the expected generalization error, and present a classical inequality in the literature (e.g., Lemma 26.2 of Shalev-Shwartz and Ben-David 2014).
Definition 3.1 (Rademacher complexity)
The empirical Rademacher complexity of a function class $\mathcal{F}(\Pi)=\{f(\pi,\cdot):\pi\in\Pi\}$ w.r.t. a dataset $(\mathbf{d}_{i})_{i=1}^{N}$ is defined as
$\mathfrak{R}\left(\mathcal{F}(\Pi)\circ(\mathbf{d}_{i})_{i=1}^{N}\right)\triangleq\ \mathbb{E}_{\sigma_{1},\ldots,\sigma_{N}}\left[\sup_{\pi\in\Pi}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\cdot f(\pi,\mathbf{d}_{i})\right],$
where $\sigma_{1},\ldots,\sigma_{n}$ are random variables independently drawn from the Rademacher distribution, i.e., ${\mathbb{P}}(\sigma_{i}=+1)={\mathbb{P}}(\sigma_{i}=-1)=1/2$.
Proposition 3.2
For any distribution $\mathcal{D}$, sample size $N$, and function class $\mathcal{F}(\Pi)=\left\{f(\pi,\cdot):\pi\in\Pi\right\}$, the following inequality holds:
[…]
The empirical Rademacher complexity measures the richness and complexity of the function class $\mathcal{F}(\Pi)$ given the dataset $(\mathbf{d}_{i})_{i=1}^{N}$. However, computing the empirical Rademacher complexity is NP-hard (see, e.g., the second paragraph on page 29 of Mohri et al. 2018). To address this challenge, combinatorial concepts, namely the VC-dimension (Vapnik and Chervonenkis 1971) and Pseudo-dimension (Pollard 1984), are introduced. Compared to Rademacher complexity, VC-dimension and Pseudo-dimension are often relatively easier to bound or estimate, providing a simpler alternative for bounding the expected generalization error. Therefore, sample complexity upper bounds can be directly inferred using the VC-dimension or Pseudo-dimension, allowing researchers to concentrate on the problem-specific task of bounding the VC-dimension or Pseudo-dimension for specific classes of algorithms of interest.
Definition 3.3 (Pseudo-dimension)
A dataset $(\mathbf{d}_{i})_{i=1}^{m}$ is shattered by a function class $\mathcal{F}(\Pi)=\{f(\pi,\cdot):\pi\in\Pi\}$ with witnesses $\tau_{1},\ldots,\tau_{m}$ if there exist $\tau_{1},\ldots,\tau_{m}\in{\mathbb{R}}$ such that for any $A\subseteq[m]$, there exists $\pi\in\Pi$ such that $f(\pi,\mathbf{d}_{i})>\tau_{i}\ \forall i\in A,$ and $f(\pi,\mathbf{d}_{i})\leq\tau_{i}\ \forall i\notin A$. The Pseudo-dimension of $\mathcal{F}(\Pi)$ denoted by $\mathrm{P\text{-}dim}(\mathcal{F}(\Pi))$ is the maximum size of a dataset that can be shattered by $\mathcal{F}(\Pi)$.
In Definition 3.3, if $f(\pi,\cdot)=\ell(\pi,\cdot)$ is the loss function, then $A$ can be interpreted as the subset of samples $i$ on which $\pi$ performs ?awfully,? i.e., it incurs a loss exceeding the desired threshold $\tau_{i}$. Note that an equivalent definition of shattering is that there exist $\tau_{1},\ldots,\tau_{m}\in{\mathbb{R}}$ such that |$\left|\{\left({\mathbb{I}}\{f(\pi,\mathbf{d}_{1})>\tau_{1}\},\ldots,{\mathbb{I}}\{f(\pi,\mathbf{d}_{m})>\tau_{m}\}\right):\pi\in\Pi\}\right|=2^{m},$ where ||$\cdot$|| ⋅ | denotes the cardinality of a set. Here, shattering means being able to realize every possible combination of binary labels on the set of points ${\mathbb{I}}\{f(\pi,\mathbf{d}_{i})>\tau_{i}\}$. Pseudo-dimension quantifies the richness of the function class to fit all possible datasets, and it generalizes the idea of VC-dimension that is designed only for binary classification. Specifically, when the function class is binary-valued, we can trivially set $\tau_{i}=0$ for all $i\in[m]$, which results in the Pseudo-dimension reducing to the VC-dimension. A higher Pseudo-dimension indicates a more complex class, which may lead to overfitting. We refer interested readers to Section 7 in the appendix, which provides some examples to illustrate the calculation of VC-dimension and Pseudo-dimension.
We next present a classical result in learning theory that uses Pseudo-dimension to derive a distribution-free bound on the Rademacher complexity (e.g., Section 5.2 of Bousquet et al. 2003).
Proposition 3.4
For any distribution $\mathcal{D}$, sample size $N$, and function class $\mathcal{F}(\Pi)=\left\{f(\pi,\cdot):\pi\in\Pi\right\}$ with $f(\pi,\cdot)$ bounded in $[0,B]$ for all $\pi\in\Pi$, there exists a positive constant $C_{0}>0$ such that
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathfrak{R}\left(\mathcal{F}(\Pi)\circ(\mathbf{d}_{i})_{i=1}^{N}\right)\right]\leq C_{0}B\sqrt{\frac{\mathrm{P\text{-}dim}(\mathcal{F}(\Pi))}{N}}.$
Combining Propositions 3.2 and 3.4, an upper bound on the Pseudo-dimension implies an upper bound on the expected generalization error. In this paper, all the three policy classes that we are interested are parameterized by real numbers. In such a case, one common technique to bound $\mathrm{P\text{-}dim}({\cal L}({\Pi}))$ from above (e.g., Balcan 2020 and Balcan et al. 2021) is to consider its dual class ${\cal L}^{*}(\Pi)\triangleq\left\{\ell(\cdot,\mathbf{d}):\mathbf{d}\in[0,U]^{T+L}\right\}$, where each function corresponds to a demand sequence $\mathbf{d}$ and is defined over the domain of $\Pi$’s parameters. We refer to $\ell(\pi,\cdot)$ and $\ell(\cdot,\mathbf{d})$ as the primal and dual loss functions, respectively. Analyzing the dual class can offer advantages over analyzing the primal class for several reasons. First, the sample space can be more complex than the parameter space, which typically lies within a Euclidean space of low dimensionality. Second, bounding the VC-dimension or Pseudo-dimension from above is challenging; it usually involves brute-forcing all possibilities when analyzing primal loss functions. By contrast, analyzing the dual loss functions provides a framework that naturally leads to an upper bound. We refer interested readers to Section 7 in the appendix for more detailed discussions. We will derive an upper bound on $\mathrm{P\text{-}dim}(\mathcal{L}(\Pi))$ by leveraging the structured nature of the dual class $\mathcal{L}^{*}(\Pi)$.
The concept of shattering can be extended to $\gamma$-shattering, introducing the term Pseudoγ-dimension (Kearns and Schapire 1994), which is also referred to as the Fat-shattering dimension in the literature. This measure, sensitive to the parameter $\gamma\geq 0$, provides a deeper understanding of the tightness of bounds on the expected generalization error. In our analysis of the policy class $\Pi_{(S^{t})}$, we will demonstrate how the Pseudoγ-dimension enables us to achieve a substantially improved upper bound, surpassing that achieved by the more traditional Pseudo-dimension approach. It is worth noting that analyzing the Pseudoγ-dimension is rare in the literature, because the Pseudo-dimension typically serves the purpose of bounding the generalization error (e.g., Morgenstern and Roughgarden 2015, Balcan et al. 2021, 2023).
Definition 3.5 (Pseudoγ-dimension)
Given $\gamma\geq 0$, a dataset $(\mathbf{d}_{i})_{i=1}^{m}$ is $\gamma$-shattered by $\mathcal{F}(\Pi)=\{f(\pi,\cdot):\pi\in\Pi\}$ with witnesses $\tau_{1},\ldots,\tau_{m}$ if there exist $\tau_{1},\ldots,\tau_{m}\in{\mathbb{R}}$ such that for any $A\subseteq[m]$, there exists $\pi\in\Pi$ such that $f(\pi,\mathbf{d}_{i})>\tau_{i}+\gamma\ \forall i\in A,$ and $f(\pi,\mathbf{d}_{i})\leq\tau_{i}-\gamma\ \forall i\notin A$. The Pseudoγ-dimension of $\mathcal{F}(\Pi)$ denoted by $\mathrm{P}_{\gamma}\text{-}\mathrm{dim}(\mathcal{F}(\Pi))$ is the maximum size of a dataset that can be $\gamma$-shattered by $\mathcal{F}(\Pi)$.
When $\gamma=0$, Pseudoγ-dimension reduces to Pseudo-dimension. We now present a useful distribution-free bound on the Rademacher complexity, using the Pseudoγ-dimension (e.g., Theorems 10 and 17 of Mendelson 2003, where the integral in our setting can be bounded between 0 and 1, because the covering number, with a range of $1$ under the $L_{2}$ norm for $\mathcal{F}(\Pi)$ that is bounded within $[-1,1]$, is 1).
Proposition 3.6
For any distribution $\mathcal{D}$, sample size $N$, and function class $\mathcal{F}(\Pi)=\left\{f(\pi,\cdot):\pi\in\Pi\right\}$ with $f(\pi,\cdot)$ bounded in $[-1,1]$ for all $\pi\in\Pi$, there exist positive constants $C_{1},C_{2}>0$ such that
[…]
4 Theoretical Results
In this section, we utilize the specific structures of loss functions that are induced by the policy classes $\Pi_{S}$, $\Pi_{(s,S)}$, and $\Pi_{(S^{t})}$ to analyze their Pseudo-dimension/Pseudoγ-dimension. By combining Propositions 3.2, 3.4 and 3.6, we derive distribution-free bounds on the expected generalization errors of these policy classes w.r.t. the horizon length $T$ and sample size $N$. This analysis also allows us to obtain the sample complexity for these policy classes.
The rest of the section is organized as follows. In Section 4.1, we summarize our technical contributions and provide an overview of the underlying intuitions and proof sketches for our results. In Section 4.2, we analyze the class of stationary base-stock policies $\Pi_{S}$, and demonstrate an upper bound of $O(\sqrt{1/N})$ for the expected generalization error of ${\cal L}(\Pi_{S})$ in Corollary 4.2. In Section 4.3, we analyze the class of $(s,S)$ policies $\Pi_{(s,S)}$, presenting an upper bound of $O(\sqrt{\log T/N})$ and a lower bound of $\Omega(\sqrt{\log T/\left(N\log\log T\right)})$ for the expected generalization error of ${\cal L}(\Pi_{(s,S)})$ in Corollaries 4.5 and 4.7, respectively. Finally, in Section 4.4, we analyze the class of non-stationary base-stock policies $\Pi_{(S^{t})}$, and show an upper bound of $O(\sqrt{1/N})$ for the expected generalization error of ${\cal L}(\Pi_{(S^{t})})$ in Corollary 4.9.
4.1 Summary of Our Results and Techniques
In this section, we provide a summary of the ideas behind our technical results in the special case with $L=0$ and $U=1$. For a horizon length $T$ and $\epsilon>0$, we state our results in terms of how many samples $N$ are required to guarantee that $\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}[\mathrm{GE}({\cal L}(\Pi))]\leq\epsilon$, under any instance defined by cost parameters $b,h,K\in[0,1]$ and distribution $\mathcal{D}$ over $[0,1]^{T}$. Such results can be converted to say that with probability at least $1-\delta$, the generalization error $R(\pi)-\hat{R}(\pi)$ of any policy $\pi\in\Pi$ is at most $\epsilon$, after adding $O(\log(1/\delta)/\epsilon^{2})$ to the number of samples $N$ (see, e.g., Lemma 26.4 and Theorem 26.5 of Shalev-Shwartz and Ben-David 2014). These results also directly bound the estimation error $R(\hat{\pi})-R(\pi^{*})$ of an empirical risk minimizer $\hat{\pi}\in\mathrm{argmin}_{\pi\in\Pi}\hat{R}(\pi)$ based on (3.1). We consider the three settings $\Pi=\Pi_{S}=[0,1]$, $\Pi=\Pi_{(s,S)}$, and $\Pi=\Pi_{(S^{t})}=[0,1]^{T}$, recalling that the policy parameter range depends on the cost parameters $b,h$ if $\Pi=\Pi_{(s,S)}$ (to be specified in Section 4.3), and that $K=0$ is assumed if $\Pi=\Pi_{(S^{t})}$.
Base-Stock (or $S$) Policies: $N=O(1/\epsilon^{2})$. We warm up by showing that the single-parameter policy class $\Pi_{S}$ indeed has sample complexity that is independent of $T$, using an argument about the dual class on the parameter space (Balcan et al. 2021), based on the convexity of the loss function w.r.t. the parameter $S$. It is also easy to see that $N=\Omega(1/\epsilon^{2})$ samples are necessary.
$(s,S)$ Policies: $N=O((\log T)/\epsilon^{2})$. We show that the Pseudo-dimension of $(s,S)$ policies is $O(\log T)$, which is sufficient. We first observe that, for a given $b,h,K$ and single demand sequence $\mathbf{d}$: (i) the minimum order quantity $\Delta\triangleq S-s$ fully determines the sequence of time periods $(t_{j})_{j\in[J]}$ at which inventory is replenished (i.e., having $y^{t}>x^{t}$), where $J$ represents the total number of reordering periods during the horizon; (ii) given a fixed sequence $(t_{j})_{j\in[J]}$, the loss function can be re-written to demonstrate that it is convex in $S$ (assuming $s$ changes to keep $\Delta$ the same). Therefore, for a fixed $(t_{j})_{j\in[J]}$, the range of $S$ that leads to a low loss on $\mathbf{d}$ is a (possibly empty) interval, and it remains to count the number of possibilities for the sequence $(t_{j})_{j\in[J]}$. A priori there would be $2^{T}$ possible subsequences of $1,\ldots,T$, but we show that $(t_{j})_{j\in[J]}$ can only change when $\Delta$ crosses over the sum of demands in an interval (i.e., crosses over $d^{t}+\cdots+d^{t^{\prime}}$ for some $1\leq t\leq t^{\prime}\leq T$). Therefore, as $\Delta$ varies, $(t_{j})_{j\in[J]}$ can change at most $O(T^{2})$ times, which means that the policy parameters achieving low loss on $\mathbf{d}$ can be defined by $O(T^{2})$ axis-aligned rectangles in the $(\Delta,S)$ plane. Ultimately we apply a counting argument about the dual class on the parameter space following Balcan et al. (2021) to argue that at most $O(\log T)$ samples can be shattered.
It may be tempting to try to prove this result by discretizing the parameter space into an $M\times M$ grid, which has $O((\log M)/\epsilon^{2})$ sample complexity. However, one can show that the approximation error from discretization remains a constant even for an arbitrary large $M$, due to the fine-tuning of $\Delta$ required based on the demand distribution; more formally,
$\sup_{\mathcal{D}}\left(\inf_{\pi=(s,S)\in\left\{\ldots,-\frac{2}{M},-\frac{1}{M},0,\frac{1}{M},\frac{2}{M},\ldots\right\}^{2}}R(\pi)\ -\inf_{\pi=(s,S)\in\Pi_{(s,S)}}R(\pi)\right)=\ \Omega(1)\ \ \ \ \ \ \forall M>0.$ (4.3)
Therefore, it is impossible to establish such a sample complexity via discretization. We defer the detailed argument of (4.3) to Section 8.1 in the appendix.
$(s,S)$ Policies: $N=\Omega((\log T/\log\log T)/\epsilon^{2})$. We show for a constant $\gamma>0$ that the Pseudoγ-dimension of $(s,S)$ policies is $\Omega(\log T/\log\log T)$, which again is sufficient due to standard reductions. Our construction involves demand sequences that start with $P-1$ 1’s followed by one $1/2$ and then all 0’s, where $P$ is a product of prime numbers. The performance of an $(s,S)$ policy depends on whether $\Delta$, which equals $S-s$ and determines the order frequency, divides $P$. The number of prime numbers that can be multiplied together without exceeding $T$ determines the number of samples that can be $\gamma$-shattered. We use the celebrated Prime Number Theorem to demonstrate that this number is $\Omega(\log T/\log\log T)$.
$(S^{t})$ Policies: $N=O(1/\epsilon^{2})$. Again, this is tight because $\Omega(1/\epsilon^{2})$ samples are necessary. We first show that the Pseudo-dimension of $(S^{t})$ policies is $\Omega(T)$ (we defer the argument to Section 8.2 in the appendix), which would suggest $N=\Omega(T/\epsilon^{2})$ and fail to provide the correct answer, so we instead analyze the stronger notion of Pseudoγ-dimension. We decompose the loss by time period, and use Talagrand’s contraction lemma (which is only possible because of the assumption that $K=0$) to show that it suffices to bound the Pseudoγ-dimension of the final inventory states $y^{T}$ induced by the policy class $\Pi_{(S^{t})}$, and show that it is $O(1/\gamma)$ for all $\gamma>0$.
We elaborate on this reduced problem. There are demand sequences $\mathbf{d}_{i}\in[0,1]^{T}$ for $i=1,\ldots,m$. An inventory policy is defined by $\pi=(S^{t})_{t=1}^{T}\in[0,1]^{T}$, and we let $y^{T}(\pi,\mathbf{d}_{i})$ denote the inventory level $y^{T}$ when policy $\pi$ is ran on demand sequence $\mathbf{d}_{i}$. The presumption is that for any subset $A\subseteq[m]$, there exists a policy $\pi$ such that inventory level $y^{T}(\pi,\mathbf{d}_{i})$ is ?high? for all $i\in A$, and ?low? for all $i\notin A$, where the thresholds for ?high? and ?low? are separated by at least $2\gamma$. We want to show that $m$ cannot be too large. Indeed, under this presumption, for each $i$ there exists a $\pi_{i}$ that ends with high inventory on $\mathbf{d}_{i}$, and low inventory on all other trajectories. We define $t_{i}$ as the last replenishment point, i.e., the last $t$ for which $S^{t}>x^{t}$, when $\pi_{i}$ is ran on $\mathbf{d}_{i}$. We essentially argue that $\mathbf{d}_{i}$ must exhibit a low total demand from $t_{i}$ to the end, compared to any other sequence $\mathbf{d}_{j}$, with the difference being at least $2\gamma$. This gives us a way of directly comparing the demand sequences $\mathbf{d}_{1},\ldots,\mathbf{d}_{m}$ without considering policies, by identifying for each $\mathbf{d}_{i}$ a time point $t_{i}$ from which its demand onward is low compared to other sequences. We link these relationships together to show that $m$ must be $O(1/\gamma)$, completing the proof.
Our result holds for arbitrary distributions $\mathcal{D}$ over $[0,1]^{T}$, but it is significant even if one assumes independence over time, i.e., $\mathcal{D}$ is a product distribution $\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{T}$. Indeed, the general theory of Guo et al. (2021) implies a sample complexity that grows linearly in the number of marginal distributions $T$ (although technically their result requires discretization or a strong monotonicity condition, we believe with some work their result can be applied here). The independent setting has also been studied in classical data-driven inventory control papers. For example, Levi et al. (2007) and Cheung and Simchi-Levi (2019) show that the number of samples required to make the estimation error $R(\hat{\pi})-R(\pi^{*})$ at most $\epsilon R(\pi^{*})$ is $O(T^{5}/\epsilon^{2})$ and $O(T^{6}/\epsilon^{2})$, respectively. Our result is not directly comparable because they are providing a multiplicative guarantee, along with other differences in the assumptions. For example, the nature of multiplicative guarantees is that one can allow for unbounded demands; however, one now needs to fix cost parameters $b,h$ before considering asymptotics in $N$ and $T$. Nonetheless, the fact that existing bounds even assuming independence grow with $T$ suggests that our $T$-independent sample complexity is quite surprising.
Beyond inventory, we believe it is interesting to encounter a setting where the stronger notion of Pseudoγ-dimension is necessary to establish the correct sample complexity. By contrast, in the data-driven algorithm design literature, Pseudo-dimension has typically been sufficient (e.g., Morgenstern and Roughgarden 2015, Balcan et al. 2021, 2023). Our problem unveils a simple reason why Pseudo-dimension may not suffice when the loss is averaged over $T$ periods, as will be explained at the beginning of Section 4.4. There, we discuss the variant involving a perishable good, whose inventory disappears at the end of each period—in which case, one period has no impact on the others, and considering the average loss over $T$ periods should be equivalent to considering a single period. Yet, the naive Pseudo-dimension is still shown to be $\Omega(T)$.
4.2 The Class of Stationary Base-Stock Policies
In this section, we analyze the class of stationary base-stock policies $\Pi_{S}=\{S:S\in[0,H]\}$, where $H$ is set to be $(L+1)U$. We note that $H$ is the maximum desirable base-stock level needed to cover the demand over $L+1$ periods. Each policy in this class is parameterized by a single base-stock level $S$. Under such a policy, if the inventory position $I^{t}$ falls below $S$ in period $t$, then an order is placed to bring the inventory position up to $S$. That is,
$q^{t}$ $=S-I^{t}=S-x^{t}-\sum_{t^{\prime}=t-L}^{t-1}q^{t^{\prime}}=\left\{\begin{array}[]{ll}S,&\text{if $t=1$};\\ d^{t-1},&\text{if $t\geq 2$}.\end{array}\right.$
It implies that $y^{t}=x^{t}+q^{t-L}=S-\sum\limits_{t^{\prime}=t-L}^{t-1}d^{t^{\prime}}$ for all $t\geq L+1$. The loss function $\ell(S,\mathbf{d})$ by the $S$ policy on demand sequence $\mathbf{d}$ is
$\ell(S,\mathbf{d})=\frac{1}{T}\sum_{t=L+1}^{T+L}\left(c\left(S-\sum_{t^{\prime}=t-L}^{t}d^{t^{\prime}}\right)+K\cdot{\mathbb{I}}\{q^{t-L}>0\}\right).$ (4.4)
Note that $\ell(S,\mathbf{d})$ is convex in $S$ for any given $\mathbf{d}$, because $c(\cdot)$ is a convex function and the fixed-cost component is independent of $S$. The following result provides an upper bound on the Pseudo-dimension of ${\cal L}(\Pi_{S})$.
Theorem 4.1 (Proved in Section 9.1)
$\mathrm{P\text{-}dim}\left({\cal L}(\Pi_{S})\right)\leq 2.$
We can now combine Proposition 3.2, Proposition 3.4 (with $B=(L+1)U+1$), and Theorem 4.1, in that order, to arrive at the following.
Corollary 4.2
For any distribution $\mathcal{D}$, sample size $N$, and horizon length $T$, the expected generalization error of $\Pi_{S}$ has the following upper bound:
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{S})\right)\right]=O\left((L+1)U\sqrt{\frac{1}{N}}\right).$
It is well-known that there exist problem instances such that the expected estimation error is lower bounded by $\Omega(\sqrt{1/N})$ (e.g., Proposition 1 of Zhang et al. 2020), which implies the tightness of the upper bound in Corollary 4.2.
Remark 4.3
According to the proof of Theorem 4.1, the conclusion that $\mathrm{P\text{-}dim}({\cal L}(\Pi_{S}))\leq 2$ essentially arises from the convexity of the loss function w.r.t. the parameter. Consequently, this argument can also be applied to upper-bound the Pseudo-dimension of other single-parameter policy classes in related inventory models. For instance, it is applicable to both the class of base-stock policies and the class of constant-order policies for the lost-sales model with lead times. In both cases, the loss function is known to be convex w.r.t. the policy parameter (e.g., Janakiraman and Roundy 2004, Xin and Goldberg 2016).
Moreover, a base-stock policy with a single $S$ is a special case of an $(s,S)$ policy when $s$ is set equal to $S$. A similar analysis can be applied to other $(s,S)$ policies with a pre-specified $S-s$ difference. For example, we can determine $S-s$ by using the following economic order quantity (EOQ) formula that allows backorders: $S-s=\sqrt{\frac{2K{\mu}(h+b)}{hb}}$ (e.g., Theorem 3.5 of Snyder and Shen 2019). Here, ${\mu}$ represents the mean demand. This enhanced version of the single-parameter policy may significantly outperform the base-stock policy numerically, as we see in Section 5.
4.3 The Class of ($s,S$) Policies
In this section, we analyze the class of ($s,S$) policies $\Pi_{(s,S)}=\left\{(s,S):\underline{H}\leq s\leq S\leq H,S\geq 0\right\}$. Under such a policy, an order is placed to bring the inventory position up to the order-up-to level $S$ when the inventory position falls below the reorder point $s$. That is,
$q^{t}$ $=(S-I^{t})\cdot{\mathbb{I}}\{I^{t}\leq s\}.$
In contrast with base-stock policies, bounding the policy parameters for $(s,S)$ policies is less clear. For example, if $h$ is small, then due to the fixed cost $K$, it may be beneficial to set $S$ to be much larger than $(L+1)U$. We bound the parameters so that the maximum holding or backlogging cost paid during any period is at most $(L+1)U$. The maximum inventory level is still $S$. Thus, we restrict $S\leq H$ where $H\triangleq(L+1)U/h$ to ensure that holding costs cannot exceed $(L+1)U$. The minimum inventory level occurs when the current inventory position $I^{t}$ is just above $s$ (a negative number), and then $(L+1)U$ demand occurs, so we pay a backlogging cost of $(-s+(L+1)U)b$. Thus, we restrict $s\geq\underline{H}$ where $\underline{H}\triangleq(L+1)U(1-1/b)\leq 0$ to ensure that backlogging costs cannot exceed $(L+1)U$. The restrictions of $s\leq S$ and $S\geq 0$ are easily seen to be without loss of optimality, resulting in the policy space $\Pi_{(s,S)}=\left\{(s,S):\underline{H}\leq s\leq S\leq H,S\geq 0\right\}$ with $\underline{H},H$ defined as above. The loss of such a policy on any demand sequence cannot exceed $(L+1)U+K\leq(L+1)U+1$.
To simplify the analysis, we focus on an equivalent definition of this policy class $\Pi_{(s,S)}=\left\{(\Delta,S):0\leq\Delta\leq H-\underline{H},0\leq S\leq H\right\}$ by defining $\Delta\triangleq S-s$. To track the inventory dynamics, we define a sequence of reordering periods $(t_{j})_{j\in[J]}$ with $1\leq t_{j}\leq t_{j+1}\leq T$, namely, $t_{j}$ is in the sequence if and only if $q^{t_{j}}>0$ and $t_{j}\leq T$. Here, $J$ represents the total number of reordering periods during the horizon, and it depends only on $\Delta$, not on $S$. We also ignore reordering periods after period $T$ because these orders will not arrive within the horizon and thus do not affect the loss. Note that $t_{1}=1$ because $x^{1}=-\infty$ by assumption, and
$t_{j+1}=\min\left\{t>t_{j}:\sum_{t^{\prime}=t_{j}}^{t-1}d^{t^{\prime}}\geq\Delta,\ t\leq T\right\}.$
Due to the lead time $L$, the loss function is
$\ell(\pi,\mathbf{d})=\frac{1}{T}\sum_{j=1}^{J}\left(K+\sum_{\hat{t}=t_{j}+L}^{t_{j+1}+L-1}c\left(S-\sum_{t^{\prime}=t_{j}}^{\hat{t}}d^{t^{\prime}}\right)\right),$
and we assume $t_{J+1}\triangleq T+1$. Based on the structure of the reordering sequence, we are able to provide an upper bound on the Pseudo-dimension of ${\cal L}(\Pi_{(s,S)})$.
Theorem 4.4 (Proved in Section 9.2)
$\mathrm{P\text{-}dim}\left({\cal L}(\Pi_{(s,S)})\right)=O(\log T)$.
As discussed earlier in Section 4.1, to our understanding, Theorem 4.4 cannot be proven by discretizing the parameter space or directly applying the result of Balcan et al. (2021). We can now combine Proposition 3.2, Proposition 3.4 (with $B=(L+1)U+1$), and Theorem 4.4, in that order, to arrive at the following.
Corollary 4.5
For any distribution $\mathcal{D}$, sample size $N$, and horizon length $T$, the expected generalization error of $\Pi_{(s,S)}$ has the following upper bound:
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(s,S)})\right)\right]=O\left((L+1)U\sqrt{\frac{\log T}{N}}\right).$
It remains to bound $\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(s,S)})\right)\right]$ from below. Note that for binary function classes, the VC-dimension provides both lower and upper bounds on the expected generalization error (e.g., Theorem 6.8 of Shalev-Shwartz and Ben-David 2014). In other words, the learnability of a binary-valued function class is equivalent to the finiteness of its VC-dimension. By contrast, a lower bound on the Pseudo-dimension of a function classes does not necessarily translate to a lower bound on the corresponding generalization error. In fact, the learnability of real-valued function classes is established through the finiteness of Pseudoγ-dimension, rather than Pseudo-dimension, and the Pseudoγ-dimension can provide both lower and upper bounds on the expected generalization error in certain contexts (e.g., Kearns and Schapire 1994, Alon et al. 1997). Therefore, we have to analyze a lower bound on the more complex Pseudoγ-dimension of ${\cal L}(\Pi_{(s,S)})$. Our next result presents such a lower bound on the Pseudoγ-dimension for some constant $\gamma>0$ by constructing an instance. Our proof involves the celebrated Prime Number Theorem and the asymptotic property of the first Chebyshev function.
Theorem 4.6 (Proved in Section 9.3)
Suppose $L\geq 0$, $U=1$, $b\in(0,1/2]$, and $h=K=0$. For any $\gamma\in[0,b/16]$, we have $\mathrm{P}_{\gamma}\text{-}\mathrm{dim}\left({\cal L}(\Pi_{(s,S)})\right)=\Omega(\log T/\log\log T)$.
We note that shattering in Theorem 4.6 is achieved using $(s,S)$ policies that satisfy the boundaries $\underline{H},H$ defined above. Note that Pseudoγ-dimension is non-increasing in $\gamma\geq 0$, implying that $\mathrm{P}_{\gamma}\text{-}\mathrm{dim}\left({\cal L}(\Pi_{(s,S)})\right)\leq\mathrm{P}_{0}\text{-}\mathrm{dim}\left({\cal L}(\Pi_{(s,S)})\right)=\mathrm{P\text{-}dim}\left({\cal L}(\Pi_{(s,S)})\right)$. Theorem 4.6 demonstrates the tightness of the order $O(\log T)$ for P-dim$\left({\cal L}(\Pi_{(s,S)})\right)$. Finally, we show that Theorem 4.6 can lead to a lower bound on the expected generalization error of $\Pi_{(s,S)}$.
Corollary 4.7 (Proved in Section 9.4)
Suppose $L\geq 0$, $U=1$, $b\in(0,1/2]$, and $h=K=0$. For any $\gamma\in(0,b/16]$ satisfying $\mathrm{P}_{\gamma}\text{-}\mathrm{dim}\left({\cal L}(\Pi_{(s,S)})\right)\geq 2$, there exists a distribution $\mathcal{D}$ such that the expected generalization error of $\Pi_{(s,S)}$ has the following lower bound for all $N\geq\mathrm{P}_{\gamma}\text{-}\mathrm{dim}\left({\cal L}(\Pi_{(s,S)})\right)$:
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(s,S)})\right)\right]=\Omega\left(\sqrt{\frac{\mathrm{P}_{\gamma}\text{-}\mathrm{dim}\left({\cal L}(\Pi_{(s,S)})\right)}{N}}\right)=\Omega\left(\sqrt{\frac{\log T}{N\log\log T}}\right).$
The proof of Corollary 4.7 uses the $\gamma$-shattered samples from Theorem 4.6 to establish a lower bound on the uniform generalization error. This shows that any error bound for $(s,S)$ policies which is established through uniform convergence must grow with $T$, unlike the case in $S$ or $(S^{t})$ policies. It is unclear to us whether an analysis of $(s,S)$ policies that eschews uniform convergence could avoid this dependence on $T$. We leave this question for future work.
4.4 The Class of Non-Stationary Base-Stock Policies
In this section, we analyze the class of base-stock policies with non-stationary base-stock levels $\Pi_{(S^{t})}=\left\{(S^{t})_{t=1}^{T+L}:(S^{t})_{t=1}^{T+L}\in[0,H]^{T+L}\right\}$ for the setting where the fixed cost is zero ($K=0$). As explained at the start of Section 4.2, we set $H=(L+1)U$. Under such a policy, the inventory position is replenished to $S^{t}$ in period $t$: $q^{t}=\max\{S^{t}-I^{t},0\}$. Recall that $I^{1}=x^{1}+\sum\limits_{t^{\prime}=1-L}^{0}q^{t^{\prime}}=-\infty$ and $y^{1}=\ldots=y^{L}=0$. It is easy to verify that
$I^{t}$ $=\ \max_{\hat{t}=1,\ldots,t-1}\left\{S^{\hat{t}}-\sum_{t^{\prime}=\hat{t}}^{t-1}d^{t^{\prime}}\right\},\ \ \forall t\geq 2,$
$y^{t}$ $=\ \max\left\{I^{t-L},S^{t-L}\right\}-\sum_{t^{\prime}=t-L}^{t-1}d^{t^{\prime}}=\ \max_{\hat{t}=1,\ldots,t-L}\left\{S^{\hat{t}}-\sum_{t^{\prime}=\hat{t}}^{t-L-1}d^{t^{\prime}}\right\}-\sum_{t^{\prime}=t-L}^{t-1}d^{t^{\prime}},\ \ \forall t\geq L+1.$
Thus, the loss in period $t\geq L+1$ is
$\ell^{t}\left((S^{t})_{t=1}^{T+L},\mathbf{d}\right)=\ c(y^{t}-d^{t})=\ c\left(\max_{\hat{t}=1,\ldots,t-L}\left\{S^{\hat{t}}-\sum_{t^{\prime}=\hat{t}}^{t-L-1}d^{t^{\prime}}\right\}-\sum_{t^{\prime}=t-L}^{t}d^{t^{\prime}}\right).$
Recall ${\cal L}(\Pi_{(S^{t})})=\left\{\ell\left((S^{t})_{t=1}^{T+L},\cdot\right):(S^{t})_{t=1}^{T+L}\in\Pi_{(S^{t})}\right\}$ is the class of loss functions induced by $\Pi_{(S^{t})}$.
As discussed in Section 4.1, the Pseudo-dimension of $(S^{t})$ policies grows linearly in $T$, leading to at best an upper bound of $O(\sqrt{T/N})$ for the expected generalization error. To provide some intuition on why this upper bound is too loose, we briefly consider a related inventory setting where the product is perishable, assuming $L=0$. In the perishable setting, leftover inventory cannot be carried over to the next period, i.e., $x^{t+1}=\min\{y^{t}-d^{t},0\}$ (instead of $x^{t+1}=y^{t}-d^{t}$). Under an $(S^{t})$ policy, inventory at time $t+1$ would always be replenished to exactly $S^{t+1}$, i.e., $y^{t+1}=\max\{S^{t+1},x^{t+1}\}=S^{t+1}$ (because $S^{t+1}\geq 0$). This means that the loss function can be decomposed as $\ell((S^{t})_{t=1}^{T},\mathbf{d})=\frac{1}{T}\sum_{t=1}^{T}c(S^{t}-d^{t})$ in the perishable setting, and we derive
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(S^{t})})\right)\right]$
$=$ $\ \mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\sup_{(S^{t})\in[0,1]^{T}}\left\{\mathbb{E}_{\mathbf{d}\sim\mathcal{D}}\left[\frac{1}{T}\sum_{t=1}^{T}c(S^{t}-d^{t})\right]-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{T}\sum_{t=1}^{T}c(S^{t}-d_{i}^{t})\right\}\right]$
$=$ $\ \frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\sup_{S^{t}\in[0,1]}\left\{\mathbb{E}_{\mathbf{d}\sim\mathcal{D}}\left[c(S^{t}-d^{t})\right]-\frac{1}{N}\sum_{i=1}^{N}c(S^{t}-d_{i}^{t})\right\}\right]$
$=$ $\ \frac{1}{T}\sum_{t=1}^{T}O(\sqrt{1/N})=O(\sqrt{1/N}),$
where the third equality comes from Corollary 4.2. This simple argument establishes an upper bound of $O(\sqrt{1/N})$ for the perishable setting, showing the Pseudo-dimension bound to be too loose. (It can be checked from the example in Section 8.2 that the Pseudo-dimension is $\Omega(T)$ even if the good is perishable.)
Of course, the simple argument above only works for the perishable setting. To establish an upper bound of $O(\sqrt{1/N})$ for the expected generalization error of $(S^{t})$ policies in the original, non-perishable setting, we now analyze the stronger notion of Pseudoγ-dimension. We use $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)$ to denote the inventory level after replenishment but before seeing demand in period $t$ under policy $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}$ on a given demand sequence $\mathbf{d}$. Note that $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)=y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)\in[-LU,H]\subseteq[-(L+1)U,(L+1)U]$ for any $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}\in\Pi_{(S^{t})}$ and $\mathbf{d}\in[0,U]^{T+L}$. We further define a class of functions representing the normalized inventory level for $t=L+1,\ldots,T+L$:
$\mathcal{Y}^{t}(\Pi_{(S^{t})})\triangleq$$\left\{y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)/((L+1)U):(S^{t^{\prime}})_{t^{\prime}=1}^{t-L}\in[0,H]^{t-L}\right\}.$
In other words, the function class $\mathcal{Y}^{t}(\Pi_{S^{t}})$ takes values in the interval $[-1,1]$.
We next provide an upper bound on the Pseudoγ-dimension of $\mathcal{Y}^{t}(\Pi_{(S^{t})})$.
Theorem 4.8 (Proved in Section 9.5)
For any $\gamma\in(0,1]$ and $t=L+1,\ldots,T+L$, $\mathrm{P}_{\gamma}\text{-}\mathrm{dim}\left(\mathcal{Y}^{t}(\Pi_{(S^{t})})\right)\leq 2/\gamma+1$.
We can now combine Proposition 3.2, Proposition 3.6, and Theorem 4.8, in that order, to arrive at the following. Note that the proof of Corollary 4.9 requires applying Talagrand’s contraction lemma to convert from analyzing losses to analyzing inventory levels, and also evaluating the integral from Proposition 3.6.
Corollary 4.9 (Proved in Section 9.6)
For any distribution $\mathcal{D}$, sample size $N$, and horizon length $T$, the expected generalization error of $\Pi_{(S^{t})}$ has the following upper bound:
$\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(S^{t})})\right)\right]=O\left((L+1)U\sqrt{\frac{1}{N}}\right).$
5 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the out-of-sample performances of the three policy classes. We use the ERM procedure to determine the output $\hat{\pi}\in\arg\inf_{\pi\in\Pi}\hat{R}(\pi)$ given the samples. Specifically, the out-of-sample (OOS) loss of policy class $\Pi$ by the ERM procedure is defined as $\text{OOS}(\Pi)\triangleq R(\hat{\pi})$. Recall that the gap between the OOS loss and the expected loss by the globally optimal policy $\pi^{\text{opt}}$ is exactly the sum of the estimation error and approximation error:
$\text{OOS}(\Pi)-R(\pi^{\text{opt}})=R(\hat{\pi})-\inf_{\pi\in\Pi}R(\pi)+\inf_{\pi\in\Pi}R(\pi)-R(\pi^{\text{opt}})=\mathrm{EE}({\cal L}(\Pi))+\mathrm{AE}({\cal L}(\Pi)).$
Our goal is to determine the best policy class among $\Pi_{S}$, $\Pi_{(s,S)}$, and $\Pi_{(S^{t})}$ based on their OOS performances. This provides insight on the tradeoff between $\mathrm{EE}({\cal L}(\Pi))$ and $\mathrm{AE}({\cal L}(\Pi))$ over different sample sizes, complementing our theoretical results that upper-bound $\mathrm{EE}({\cal L}(\Pi))$.
In the rest of the section, we numerically evaluate the expected OOS loss w.r.t. the sample size $N$. Specifically, in Section 5.1, we focus on $\Pi_{(s,S)}$ and $\Pi_{S}$ when $K>0$. In Section 5.2, our evaluation concerns $\Pi_{(S^{t})}$ and $\Pi_{S}$ when $K=0$. We also demonstrate the benefits of integrating inventory theory into black-box learning machines in Section 5.3 by comparing two approaches: one involves learning a base-stock policy by leveraging inventory theory, and the other focuses on learning order quantities without relying on inventory theory. We select $N\in\{2,5,10,20,50,100\}$, and set $L=0$, $h=1$ and $b=9$, which corresponds to the service level $\frac{b}{h+b}=90\%$. Note that we do not normalize $b$ and $K$ to lie within the range $[0,1]$ in our numerical experiments. Because $\mathrm{EE}({\cal L}(\Pi))$ decreases as $N$ increases, $R(\pi^{*})$ can be estimated by using $\text{OOS}(\Pi)$ with a large $N$, and we set $N=300$ in simulations. To approximate $\text{OOS}(\Pi)=R(\hat{\pi})=\mathbb{E}_{\mathbf{d}\sim\mathcal{D}}[\ell(\hat{\pi},\mathbf{d})]$, we calculate the average of $\ell(\hat{\pi},\mathbf{d}_{j})$ for 3,000 randomly generated samples $(\mathbf{d}_{j})_{j}$. To evaluate the expected OOS loss w.r.t. the training dataset $\mathbf{d}_{1},\ldots,\mathbf{d}_{N}$, we repeat the experiments 100 times and record the average value of $\text{OOS}(\Pi)$. In our experiments, we do not restrict the ranges of policy parameters using constants such as $H$ and $\underline{H}$; however, we do restrict all demand realizations and order quantities to integer values.
5.1 $\Pi_{(s,S)}$ vs. $\Pi_{S}$ when $K>0$
In this section, we compare the expected OOS losses of $\Pi_{(s,S)}$ and $\Pi_{S}$ when $K>0$. In our experiments, we set $T=30$ and investigate demand sequences that are i.i.d. and drawn from a normal distribution $N(\mu,\sigma^{2})$ with $\mu=9$ and $\sigma=4.5$. These sequences are then truncated and rounded into non-negative integers. We select different values of $K$ based on the refined EOQ formula $\Delta=\sqrt{\frac{2K\mu(h+b)}{hb}}$ (as mentioned in Section 4.2). More specifically, we set $K(P)=\frac{P^{2}\mu hb}{2(h+b)}=4.05P^{2}$ as the value of $K$, so that replenishment occurs approximately every $P$ periods, following from $\Delta/\mu=P$. Recall that $\Pi_{(s,S)}$ is asymptotically optimal (as $T\to\infty$) under i.i.d. demand distributions. Instead of plotting the expected OOS losses, we plot the ratio of the expected OOS losses of $\Pi_{(s,S)}$ and $\Pi_{S}$ relative to $R(\pi^{*})$, because ratios are easier to interpret. To capture the empirical policies learned from the training dataset, we calculate the average of the empirical policies obtained across $100$ repeated experiments, each using $N=300$ samples.
We present the results for $\Pi_{(s,S)}$ and $\Pi_{S}$ with different values of $K$ in Figure 2. The difference between the dashed lines represents a (normalized) approximation error for $\Pi_{S}$, since $(s,S)$ policies are near-optimal. We observe that the approximation error of $\Pi_{S}$ increases as $K$ increases. This is because a larger $K$ leads to less frequent replenishments under $(s,S)$ policies, whereas $S$ policies keep replenishing almost every period, incurring high fixed costs. It further suggests that comparing $\text{OOS}(\Pi_{S})$ and $\text{OOS}(\Pi_{(s,S)})$ is meaningless, which motivates us to consider a variant of $(s,S)$ policies. This variant would have a pre-specified $S-s$ difference, determined by the refined EOQ formula.
(a) $K(1)=4.05\times 1^{2}=4.05$
(b) $K(2)=4.05\times 2^{2}=16.2$
Figure 2: The expected OOS loss ratios of $\Pi_{(s,S)}$ and $\Pi_{S}$ when $K>0$ and demands are i.i.d. and drawn from a normal distribution $N(9,4.5^{2})$. The averaged empirical policies learned from $N=300$ samples across 100 experiments are: (a) $\widehat{({s},{S})}=(11.43,15.02)$ and $\hat{S}=15.0$; (b) $\widehat{({s},{S})}=(8.96,23.38)$ and $\hat{S}=15.0$.
Particularly, let $\bar{\mu}$ denote the empirical mean of the training demand samples, and define $\bar{\Delta}\triangleq\sqrt{\frac{2K\bar{\mu}(h+b)}{hb}}$. We consider this policy class defined as $\Pi_{\text{EOQ}}\triangleq\left\{(s,S):S\in[0,H],s=S-[\bar{\Delta}]\right\}$, where $[x]$ denote the nearest integer to $x$. Note that the demand is i.i.d. drawn from $N(9,4.5^{2})$, and $\bar{\mu}$ is the sample mean of $N\times T$ i.i.d. demand values. In addition, $\Pi_{\text{EOQ}}$ is expected to have much smaller approximation errors than $\Pi_{S}$ due to the fixed costs. Therefore, $\Pi_{\text{EOQ}}$ can be regarded as an enhanced version of $\Pi_{S}$.
(a) $K(2)=4.05\times 2^{2}=16.2$
(b) $K(3)=4.05\times 3^{2}=36.45$
(c) $K(5)=4.05\times 5^{2}=101.25$
Figure 3: The expected OOS loss ratios of $\Pi_{(s,S)}$ and $\Pi_{\text{EOQ}}$ when $K>0$ and demands are i.i.d. and drawn from a normal distribution $N(9,4.5^{2})$. The averaged empirical policies learned from $N=300$ samples across 100 experiments are reported as follow: (a) $\widehat{({s},{S})}=(8.99,26.02)$ and $\hat{S}=26.02$, $[\bar{\Delta}]=18$; (b) $\widehat{({s},{S})}=(7.36,29.98)$ and $\hat{S}=33.96$, $[\bar{\Delta}]=27$; (c) $\widehat{({s},{S})}=(4.91,44.72)$ and $\hat{S}=49.02$, $[\bar{\Delta}]=45$.
We present the results for $\Pi_{(s,S)}$ and $\Pi_{\text{EOQ}}$ with different values of $K$ in Figure 3. First, we are interested in the estimation errors of $\Pi_{(s,S)}$ and $\Pi_{\text{EOQ}}$. Because the curves represent the expected OOS losses for the policy classes, and the dotted horizontal lines correspond to the simulated values of $R(\pi^{*})$, the gap between the OOS curve and the dotted line represents a (normalized) estimation error. Similar to Figure 2, all parts in Figure 3 show that $(s,S)$ policies have much larger estimation errors than the single-parameter policies, confirming that learning ${(s,S)}$ policies is more challenging. Moreover, we are particularly interested in the intersection points of the OOS curves, with a special focus on their relation to sample size. This is because the intersection point represents the sample size at which one would want to switch from the simpler policy class $\Pi_{\text{EOQ}}$ (or $\Pi_{S}$) to the richer policy class $\Pi_{(s,S)}$. As illustrated in these figures, even with the simplest i.i.d. demand distribution, the intersection point is noticeably non-zero due to the large estimation error of $\Pi_{(s,S)}$. It demonstrates that opting for the simpler policy class $\Pi_{\text{EOQ}}$ is preferable to opting for $\Pi_{(s,S)}$, which aligns with the principle ?learning less is more.? For larger values of $K$, both the empirical gap $\hat{\Delta}=\hat{S}-\hat{s}$ observed in the $(s,S)$ policies and the gap $[\bar{\Delta}]$ observed in the EOQ policies increase. However, the difference between these two gaps, $[\bar{\Delta}]-\hat{\Delta}$, remains relatively stable. As a result, as $K$ increases, the fixed cost component dominates the other two cost components, and the total number of replenishments under $\Pi_{\text{EOQ}}$ becomes close to that under $\Pi_{(s,S)}$, leading to a relatively smaller difference in approximation error. This is seen in Figure 3, which shows that the gap between the two dotted lines narrows, and the intersection point shifts rightward as $K$ increases.
5.2 $\Pi_{(S^{t})}$ vs. $\Pi_{S}$ when $K=0$
In this section, we compare the expected OOS losses of $\Pi_{S}$ and $\Pi_{(S^{t})}$ when $K=0$. We set $T=10$ and investigate demand sequences $\mathbf{d}=(d^{t})_{t=1}^{T}$ that are independently drawn from normal distributions with non-stationary means and variances across time, namely, $d^{t}\sim N(\mu^{t},\sigma^{t})$ $\forall t\in[T]$. These sequences are then truncated and rounded into non-negative integers. Here, $\mu^{t}$ and $\sigma^{t}$ represent the mean and standard deviation of the demand at period $t$, respectively, and we set $\sigma^{t}=\mu^{t}/2$ for all $t\in[T]$. We consider two types of demand sequences: almost stationary demand and highly non-stationary demand. Specifically, for the almost stationary demand, $\mu^{t}$ is i.i.d. and drawn from a uniform distribution $U(7,11)$. For the highly non-stationary demand, $\mu^{t}$ is i.i.d. and drawn from a uniform distribution $U(1,17)$. In both settings, we generate 10 sets of mean vectors and record the average value of $\text{OOS}(\Pi)$. Because $\Pi_{(S^{t})}$ is optimal under independent and non-stationary demand distributions, we plot the ratio of the expected OOS losses of $\Pi_{S}$ and $\Pi_{(S^{t})}$ relative to $R(\pi^{*})$, estimated using $N=300$ samples.
(a) Almost Stationary
(b) Highly Non-Stationary
Figure 4: The expected OOS loss ratios of $\Pi_{(S^{t})}$ and $\Pi_{S}$ when $K=0$.
We present the results for the two types of demand distributions in Figure 4. First, it is shown that $\mathrm{EE}({\cal L}(\Pi_{(S^{t})}))$ exceeds $\mathrm{EE}({\cal L}(\Pi_{S}))$, confirming that learning multiple parameters can be more challenging than learning a single parameter. However, as $N$ increases, the estimation error of $\Pi_{(S^{t})}$ decreases very quickly, much quicker than that of $\Pi_{(s,S)}$ as shown in Figure 3. This complements our theoretical finding that $(S^{t})$ policies have a better generalization bound and lower estimation error than $(s,S)$ policies.
In the setting with almost stationary demand, the demand distribution closely resembles that of the i.i.d. setting, implying that single base-stock policies can effectively approximate the optimal $(S^{t})$ policy. On the other hand, when demand becomes highly non-stationary, $S$ policies have a large approximation error that dominates the estimation error of $(S^{t})$ policies. In fact, the intersection point vanishes (i.e., becomes zero), making it always better to learn the richer class of $(S^{t})$ policies. Note that Figure 4(a) is designed to closely approximate stationary conditions for illustrative purposes; however, typical scenarios are expected to resemble more closely the situation depicted in Figure 4(b). This suggests that learning $\Pi_{(S^{t})}$ should generally be preferred over learning $\Pi_{S}$, even at small sample sizes, because $\Pi_{(S^{t})}$ is quite safe from overfitting, complementing our theoretical findings.
5.3 Base-Stock vs. Order Quantity
Inventory theory has evolved over more than 70 years, beginning with the pioneering work of Arrow et al. (1951). A cornerstone of this field is the demonstration of the optimality of a base-stock policy for backlogged models when $K=0$, a fundamental result in inventory theory. This knowledge has naturally led to the practice of directly learning base-stock policies based on historical demand sequences. Without this foundational understanding, one might instead learn a complex function mapping each inventory state to a corresponding order quantity. The objective of this section is to highlight the critical role that domain knowledge plays in the field of inventory theory. We support our argument by comparing the outcomes of learning a base-stock policy with the advantage of inventory theory versus learning order quantities in its absence. We refer to these two approaches as ?learning-base-stock? and ?learning-order-quantity,? respectively.
Let us set $K=0$, $T=30$ and $L=1$. We consider demand sequences that are i.i.d. and drawn from a discrete uniform distribution $U\{0,\ldots,\bar{D}\}$, where $d^{t}=0,1,\ldots,\bar{D}$ with equal probability. We test cases where $\bar{D}=2$ and $\bar{D}=3$. In these settings, learning order quantities is equivalent to learning a function $f(\cdot)$ that maps the inventory level $y^{t}$ after replenishment at the time $t$ to an order quantity $q^{t}$. We plot the ratio of the expected OOS losses to the OOS loss of learning order quantities using $N=300$ samples. We investigate both backlogged and lost-sales models: the former, which is the main focus of this paper, is well known for the optimality of a base-stock policy; the latter is notoriously challenging to analyze within the inventory literature (e.g., Zipkin 2008, Xin and Goldberg 2016, Xin 2021), making it particularly interesting to illustrate the power of domain knowledge in inventory management.
We first compare the two approaches in the backlogged model. Note that the state space can be condensed to $y^{t}\in\{-\bar{D},\ldots,0,\ldots,2\bar{D}\}$, given that $d^{t}\in\{0,1,\ldots,\bar{D}\}$. Consequently, there are $13\bar{D}+$ 1 parameters to learn. In addition, the range of $f(y^{t})$ can be narrowed to $\left\{(-y^{t})^{+},\ldots,2\bar{D}-y^{t}\right\}$. This reduction implicitly incorporates some domain knowledge about inventory dynamics. The results are presented in Figure 5. The estimation error associated with learning a single base-stock level is significantly lower than that of learning $13\bar{D}+$ 1 parameters. In addition, given that an optimal base-stock policy exists, both approaches yield zero approximation error. As a consequence, the learning-base-stock approach dominates the learning-order-quantity approach for all sample sizes. Furthermore, their performance gap widens for larger values of $\bar{D}$, when the learning-order-quantity approach faces an increased number of parameters to learn.
(a) $d^{t}\sim U\{0,1,2\}$
(b) $d^{t}\sim U\{0,1,2,3\}$
Figure 5: The expected OOS loss ratios of learning-base-stock and learning-order-quantity in the backlogged model.
Next, we conduct a similar comparison in the more complicated lost-sales model with lead times. The lost-sales assumption is especially appropriate in online retail settings, where customers can easily switch among different competing retailers. In this lost-sales model, the optimal policy often involves complex, state-dependent decisions and is challenging to compute due to the curse of dimensionality. Although base-stock policies may not always be optimal, inventory theory suggests that they become asymptotically optimal as shortage costs increase (Huh et al. 2009, Wei et al. 2021). These policies are also expected to perform well for short lead times. Similarly, the state space can be condensed to $y^{t}\in\{0,1,\ldots,2\bar{D}\}$ (i.e., $12\bar{D}+$ 1 parameters to learn), and the range of $f(y^{t})$ can be narrowed to $f(y^{t})\in\{0,1,\ldots,2\bar{D}-y^{t}\}$. The results are presented in Figure 6. Because the loss function is also convex w.r.t. the base-stock level in the lost-sales model (e.g., Janakiraman and Roundy 2004), the estimation error for the learning-base-stock approach is also $O(\sqrt{1/N})$, as discussed in Remark 4.3. This is also evidenced in Figure 6, where it is shown that the estimation error for the learning-base-stock approach is significantly lower than that for the learning-order-quantity approach. In addition, the intersection point between the two approaches is relatively large, suggesting that the simpler learning-base-stock approach is preferable for small sample sizes. This observation aligns with the principle that ?learning less is more.? Furthermore, as $\bar{D}$ increases, the intersection point shifts rightward, even though the approximation error (distance between the dotted lines) has grown. This suggests that overfitting becomes a much more dominant issue as the problem scale increases. We note that the problem size is quite small in the settings studied here, with $L=1$ and discrete demand $d^{t}\in\{0,\ldots,\bar{D}\}$, involving only $12\bar{D}+$ 1 states. The size of the state space can be significantly larger for unbounded distributions and also increases exponentially with $L$. For example, as illustrated in Table 2 of Zipkin (2008), with a Geometric demand distribution and the same cost parameters ($h=1$ and $b=9$), an increase in $L$ from 1 to 4 results in the size of the effective state space—after applying a state-reduction technique—growing from 21 to 52,722. In such settings with large state spaces, we can expect the intersection point to be significantly shifted to the right, implying that learning-base-stock outperforms learning-order-quantity across a wide range of sample sizes. This highlights the importance of focusing on well-structured policy classes when data is limited, emphasizing the value of domain knowledge derived from 70 years of inventory theory. Finally, we note that the idea of learning well-structured policies is not unprecedented in the online learning literature, particularly within the context of the lost-sales inventory model with lead times and censored demand. Notably, all existing studies focus on learning the optimal base-stock (or other well-structured) policy, which may not necessarily be optimal (e.g., Zhang et al. 2020, Lyu et al. 2024).
(a) $d^{t}\in\{0,1,2\}$
(b) $d^{t}\in\{0,1,2,3\}$
Figure 6: The expected OOS loss ratios of learning-base-stock and learning-order-quantity in the lost-sales model.
6 Conclusion
This paper adopts a novel approach by analyzing generalization guarantees of several well-known classes of inventory policies, leveraging the celebrated VC theory. Both our theory and simulations reveal important managerial insights into the principle of ?learning less is more? in inventory management. Depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies. Moreover, the number of parameters in a policy class may not accurately measure overfitting. For example, the generalization error for $(S^{t})$ policies is comparable to that of $(s,S)$ policies, despite having significantly more parameters.
Our work has the potential for broader, significant impacts. In recent years, many studies have applied state-of-the-art deep learning approaches to solve large-scale inventory problems. We believe our structured approach of learning base-stock levels instead of order quantities can potentially enhance existing deep learning techniques, which is particularly valuable in retailing where relevant data is often limited. In other words, it is important to leverage the domain knowledge from 70 years of inventory theory. The idea of leveraging academic knowledge and restricting to policy structures that have been well-established by theory can potentially translate to other problems too.
[…]
The authors thank Matias Alvo for providing useful comments on an early draft.

### v2

@@
 
-[…]
 […]
@@
 In particular, we prove generalization guarantees for learning several well-known classes of inventory policies, including base-stock and $(s,S)$ policies, by leveraging the celebrated Vapnik-Chervonenkis (VC) theory.
-We apply the concepts of the Pseudo-dimension and Fat-shattering dimension from VC theory to determine the generalizability of inventory policies, that is, the difference between an inventory policy’s performance on training data and its expected performance on unseen data.
+We apply the concepts of the Pseudo-dimension and Fat-shattering dimension from VC theory to determine the generalization error of inventory policies, that is, the difference between an inventory policy’s performance on training data and its expected performance on unseen data.
 We focus on a classical setting without contexts, but allow for an arbitrary distribution over demand sequences and do not make any assumptions such as independence over time.
@@
 in inventory management: depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies to minimize overfitting errors.
-Second, the number of parameters in a policy class may not be the correct measure of overfitting error: in fact, the class of policies defined by $T$ time-varying base-stock levels exhibits a generalization error comparable to that of the two-parameter $(s,S)$ policy class.
+Second, the number of parameters in a policy class may not be the correct measure of overfitting error: in fact, the class of policies defined by $T$ time-varying base-stock levels exhibits a generalization error an order of magnitude lower than that of the two-parameter $(s,S)$ policy class.
 Finally, our research suggests situations in which it could be beneficial to incorporate the concepts of base-stock and inventory position into black-box learning machines, instead of having these machines directly learn the order quantity actions.
@@
 […]
-May 1, 2024
+July 7, 2024
 But it is the VC dimension rather than the number of parameters of the set of functions that defines the generalization ability of a learning machine.(Vapnik 1998, Page 160)
@@
 2022).
-The latter papers highlight the power of black-box neural networks, which do not require explicit knowledge about the structure of optimal inventory policies to achieve high performance.
+The latter papers highlight the power of black-box neural networks, which do not require explicit knowledge of the structure of optimal inventory policies to achieve high performance.
 By contrast, our paper demonstrates the potential benefits of adhering to policy structures that have been well-established in 70 years of inventory theory, beginning with Arrow et al.
 (1951).
-We derive learning theory results and complement them with simulations.
-We analyze the fundamental classes of base-stock and $(s,S)$ policies, which are optimal in various settings with independent demand over time (e.g., Snyder and Shen 2019).
-We approach this as a supervised learning problem over demand sequences, in the spirit of Madeka et al.
-(2022), Alvo et al.
-(2023), but focus on a classical setting without contexts.
-We use the celebrated Vapnik-Chervonenkis (VC) theory to bound the learning-theoretic complexity of these policy classes when trained offline on past demand sequences, without assuming they are independent over time as this is often impossible to verify in practice.
-We provide statistical learning guarantees for this training procedure, which resembles the ?hindsight?
-reinforcement learning approaches mentioned earlier.
+Instead of learning arbitrary ordering actions, we learn inventory policies that are defined by the well-established concepts of base-stock levels and reorder points.
+Not only does this further improve empirical performance, but it is also founded on theoretical guarantees for learning the best policies within these classes.
+In particular, we analyze the fundamental classes of base-stock and $(s,S)$ policies, which are optimal in various settings with independent demand over time (e.g., Simchi-Levi et al.
+2014, Snyder and Shen 2019).
+We focus on a non-contextual setting where we are given samples from an unknown underlying distribution of demand sequences $(d^{t})_{t=1}^{T}$ (whose entries can be arbitrarily correlated over time $t$) over a $T$-period horizon.
+We view this parsimonious, context-free setting as an important starting point for understanding the performance of the supervised learning approach described by Madeka et al.
+(2022) and Alvo et al.
+(2023).
+This approach still captures the high-level goal of incorporating arbitrary demand patterns in a model-free way.
+We analyze the estimation error, which is the performance gap between the empirical policy learned from training data and the best-in-class policy when evaluated on the true distribution.
+The estimation error depends on the number of samples, which is the general focus of analysis in data-driven Newsvendor and inventory problems (see Levi et al.
+2007).
+However, papers in this literature assume that the random demands $d^{t}$ are independent over time $t$, and the estimation error is bounded by analyzing the dynamic program for independent demands.
+Our analysis completely relaxes the independence assumption, motivated by the aforementioned goal of supervised learning over correlated sequences, and in some cases even improves the guarantee on estimation error.
+Although seemingly paradoxical, eliminating the independence assumption necessitates a completely different analysis based on the celebrated Vapnik-Chervonenkis (VC) theory, which proves useful even for independent demands.
 VC theory is a foundational framework in the field of statistical learning theory (Vapnik and Chervonenkis 1971).
@@
 These dimensions can be used to provide a distribution-free upper bound on the Rademacher complexity, and overall, quantify the capability of a hypothesis class to learn complex patterns.
-A hypothesis class with a higher dimension generally possesses a greater potential to learn a more complex pattern from data, but may suffer from reduced generalizability due to overfitting, leading to the need for a larger training dataset to ensure good performance of the learned hypothesis.
+A hypothesis class with a higher dimension generally possesses a greater potential to learn more complex patterns from data.
+However, it may suffer from reduced generalizability due to overfitting, necessitating a larger training dataset to ensure good performance of the learned hypothesis.
 In the context of inventory management, VC theory can be a powerful tool for establishing rigorous bounds on the generalization error of inventory policy classes, namely, the difference between an inventory policy’s performance on training data and its expected performance on true data.
 This is useful because, as we will show, naive metrics such as the number of parameters do not necessarily indicate generalizability.
+Bounds on the generalization error can then be directly translated back into bounds on the estimation error of learning best-in-class policies.
 We instead use VC theory to indicate generalizability, and corroborate our theoretical results using simulations.
-All in all, this reveals how effective each of these policy class restrictions is in terms of reducing overfitting, complementing the large body of inventory theory that already characterizes when each of these policy classes are optimal.
+All in all, this reveals how effective each of these policy class restrictions is in terms of reducing overfitting, complementing the large body of inventory theory that already characterizes when each of these policy classes is optimal.
 1.1 ?Learning Less is More? Phenomenon
@@
 Demands, cost parameters, and policy parameters are normalized so that given any $T$, the performance of any policy on any demand realization is bounded by a constant.
-The demand realization indicates the demand for the good during each time period, drawn from an arbitrary distribution over sequences of length $T$ (that could exhibit correlations over time).
+The realized demand vector $(d^{t})_{t=1}^{T}$ indicates the demand for the good during each time period, drawn from an arbitrary distribution over sequences of length $T$.
 We assume access to $N$ independent and identically distributed (i.i.d.) samples of sequences drawn from the same distribution, and investigate how the uniform generalization error for a policy class $\Pi$ (difference between expected performance on the true distribution, and average empirical performance over the $N$ samples, taking the supremum over policies in $\Pi$) scales with $N$ and $T$.
@@
 setting.
+More recently, Qin et al.
+(2023) have derived results for the same additive setting as ours, and even under independence their sample complexity for $(S^{t})$ policies grows with $T$.
 Therefore, we believe the non-dependence on $T$ in our additive setting to be quite surprising.
@@
 Levi et al.
-(2007) and Cheung and Simchi-Levi (2019) both consider dynamic models and prove sample complexity bounds on base-stock policies under the assumption that demands are independent across time but not necessarily identically distributed.
+(2007), Cheung and Simchi-Levi (2019), and Qin et al.
+(2023) all consider dynamic models and prove sample complexity bounds on base-stock policies under the assumption that demands are independent across time but not necessarily identically distributed.
 Huh and Rusmevichientong (2009) use stochastic gradient descent to learn inventory policies for demand that is both independent and identically distributed over time.
@@
 (2021) provide an upper bound on the regret of learning the optimal $(s,S)$ policy.
+Fan et al.
+(2024) demonstrate that one can beat ?best-possible?
+bounds derived for general MDPs by leveraging the structural properties of inventory dynamics under $(s,S)$ policies.
 It is quite surprising that the study of VC-type complexity, even for the simplest classes such as base-stock and ($s,S$) policies, is missing; this is the gap we aim to fill with this paper.
-It is worth noting that Han et al.
-(2023) apply existing bounds on the Pseudo-dimension of neural networks directly to solve a Newsvendor problem.
+It is worth noting that, recently, Chen et al.
+(2022) and Han et al.
+(2023) applied existing bounds on the generalization error of neural networks directly to a Newsvendor problem.
 Their model is based on the contextual Newsvendor model that was pioneered by Ban and Rudin (2019).
@@
 Here, the $N$ samples can be interpreted as the demands of $N$ similar products over $(T+L)$ periods, or as $N$ records of one product’s demands over $(T+L)$ consecutive periods (e.g., each sample $i=1,\ldots,N$ represents one year and consists of $T+L$ periods).
+The samples can also correspond to the life cycles of the $N$ past generations of a product (e.g.
+a magazine or video game), whose sales are correlated within cycle (e.g.
+one generation was popular and had high demands throughout its life cycle) but each cycle can be viewed as an i.i.d.
+realization.
 3.1 Estimation Error
@@
 $\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}[\mathrm{EE}({\cal L}(\Pi))]$ $=\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[R(\hat{\pi})-\hat{R}(\hat{\pi})+\hat{R}(\hat{\pi})-\hat{R}(\pi^{*})+\hat{R}(\pi^{*})-R(\pi^{*})\right]$
+$\leq\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[R(\hat{\pi})-\hat{R}(\hat{\pi})\right]$
 $\leq\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\sup_{\pi\in\Pi}\left\{R({\pi})-\hat{R}({\pi})\right\}\right],$ (3.1)
-where the inequality comes from the optimality of $\hat{\pi}$ and the fact that $\mathbb{E}\left[\hat{R}(\pi^{*})-R(\pi^{*})\right]=0$.
-Note that from an upper bound of the expected estimation error, one can directly derive the sample complexity, i.e., the number of samples required to guarantee a given estimation error with a given high probability.
+where the first inequality comes from the optimality of $\hat{\pi}$ and the fact that $\mathbb{E}\left[\hat{R}(\pi^{*})-R(\pi^{*})\right]=0$.
 The supremum term in (3.1) is referred to as the (uniform) generalization error (GE) or the representativeness of the policy class $\Pi$ in the literature, which describes the worst-case scenario of overfitting within the policy class.
@@
 $\mathrm{GE}({\cal L}(\Pi))\triangleq\sup_{\pi\in\Pi}\left\{R({\pi})-\hat{R}({\pi})\right\}.$ (3.2)
-It quantifies the generalization ability of an algorithm (or a policy), namely, the gap between the algorithm’s in-sample performance on the training set and its out-of-sample performance.
-Note that another widely accepted definition of uniform generalization error is $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$, which is commonly used for deriving high-probability sample complexity bounds.
-The analysis based on this definition is analogous to that by (3.2), expect for a constant factor.
+It quantifies the generalization ability of a policy class $\Pi$, namely, the worst-case gap between in-sample and out-of-sample performances of a policy within $\Pi$.
 In this paper, our primary emphasis is on analyzing the expected generalization error w.r.t.
@@
 $T$ varies among different policy classes, by leveraging the concepts of VC-Dimension, Pseudo-Dimension and Pseudoγ-Dimension.
+We remark that the literature often focuses on high-probability bounds for the estimation error instead of expectation bounds, in which case the definition of generalization error can be $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$.
+Regardless, our expectation upper bounds imply high-probability upper bounds after applying standard tools such as McDiarmid’s inequality (e.g., Theorem 26.5 of Shalev-Shwartz and Ben-David 2014), and our analysis of $\sup\limits_{\pi\in\Pi}\left\{R({\pi})-\hat{R}({\pi})\right\}$ implies upper bounds on $\sup\limits_{\pi\in\Pi}\left\{\left|R({\pi})-\hat{R}({\pi})\right|\right\}$ after scaling by a constant factor.
 3.2 Approximation Error
@@
 These three classes of policies have been extensively studied in the inventory literature and are widely implemented in practice.
-We next review several classical results for finite-horizon backlogged inventory models in the stochastic inventory control literature (e.g., Snyder and Shen 2019), demonstrating the conditions under which these policies achieve optimality (i.e., zero approximation error).
+We next review several classical results for finite-horizon backlogged inventory models in the stochastic inventory control literature (e.g., Simchi-Levi et al.
+2014, Snyder and Shen 2019), demonstrating the conditions under which these policies achieve optimality (i.e., zero approximation error).
   • In settings where the fixed cost $K=0$ and demands are i.i.d.
 across time, an optimal base-stock policy exists with a stationary base-stock level for each period, falling within the class $\Pi_{S}$ with $H=(L+1)U$.
-  • In settings where $K>0$ and demands are i.i.d.
-across time, an $(s,S)$ policy with both stationary order-up-to level $S$ and reorder point $s$ is asymptotically optimal (as $T\to\infty$), falling within the class $\Pi_{(s,S)}$ where $H$ and $\underline{H}$ are selected appropriately based on the cost parameters (to be further discussed in Section 4.3).
+  • In settings where $K>0$ and demands are i.i.d. across time, an $(s,S)$ policy with both stationary order-up-to level $S$ and reorder point $s$ is asymptotically optimal (as $T\to\infty$).
   • In settings where the fixed cost $K=0$ and demands are independently but not necessarily identically distributed across time, an optimal base-stock policy exists.
@@
 Compared to Rademacher complexity, VC-dimension and Pseudo-dimension are often relatively easier to bound or estimate, providing a simpler alternative for bounding the expected generalization error.
-Therefore, sample complexity upper bounds can be directly inferred using the VC-dimension or Pseudo-dimension, allowing researchers to concentrate on the problem-specific task of bounding the VC-dimension or Pseudo-dimension for specific classes of algorithms of interest.
+Therefore, sample complexity upper bounds can be directly inferred using the VC-dimension or Pseudo-dimension, allowing researchers to concentrate on the problem-specific task of bounding the VC-dimension or Pseudo-dimension for specific policy classes of interest.
 Definition 3.3 (Pseudo-dimension)
@@
 Nonetheless, the fact that existing bounds even assuming independence grow with $T$ suggests that our $T$-independent sample complexity is quite surprising.
+The surprisingness of our result is further echoed in a recent paper by Qin et al.
+(2023), who study the same additive setting as ours.
+Their sample complexity of $(S^{t})$ policies is $O(T/\epsilon^{2})$, also growing with $T$.
+(Their stated sample complexity is $O(T^{3}/\epsilon^{2})$ because their cost function does not divide by $T$.)
+Our $T$-independent result also relates to a broad literature on studying dimension-dependence sample complexity bounds in stochastic programming and learning theory.
+For example, in convex stochastic programming, Liu and Tong (2024) present a dimension-independent sample complexity bound for sample average approximation under a set of assumptions, including strong convexity.
+In contrast, the loss function of $(S^{t})$ policies in our setting is not even convex in $(S^{t})$ (e.g., Kunnumkal and Topaloglu 2008).
+For generic Lipschitz-continuous loss functions, Shalev-Shwartz et al.
+(2009) establish a sample complexity bound that depends on the dimension of the decision variable (i.e., $T$ in our setting).
+However, we demonstrate that in our specific inventory context, the $T$-dependence is unnecessary.
+Considering all these aspects, we find our $T$-independent result to be quite surprising.
 Beyond inventory, we believe it is interesting to encounter a setting where the stronger notion of Pseudoγ-dimension is necessary to establish the correct sample complexity.
@@
 In other words, the learnability of a binary-valued function class is equivalent to the finiteness of its VC-dimension.
-By contrast, a lower bound on the Pseudo-dimension of a function classes does not necessarily translate to a lower bound on the corresponding generalization error.
+By contrast, a lower bound on the Pseudo-dimension of real-valued function classes does not necessarily translate to a lower bound on the corresponding generalization error.
 In fact, the learnability of real-valued function classes is established through the finiteness of Pseudoγ-dimension, rather than Pseudo-dimension, and the Pseudoγ-dimension can provide both lower and upper bounds on the expected generalization error in certain contexts (e.g., Kearns and Schapire 1994, Alon et al.
@@
 We use $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)$ to denote the inventory level after replenishment but before seeing demand in period $t$ under policy $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}$ on a given demand sequence $\mathbf{d}$.
-Note that $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)=y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)\in[-LU,H]\subseteq[-(L+1)U,(L+1)U]$ for any $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}\in\Pi_{(S^{t})}$ and $\mathbf{d}\in[0,U]^{T+L}$.
+Note that $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)=y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)\in[-LU,(L+1)U]\subseteq[-(L+1)U,(L+1)U]$ for any $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}\in\Pi_{(S^{t})}$ and $\mathbf{d}\in[0,U]^{T+L}$.
 We further define a class of functions representing the normalized inventory level for $t=L+1,\ldots,T+L$:
-$\mathcal{Y}^{t}(\Pi_{(S^{t})})\triangleq$$\left\{y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)/((L+1)U):(S^{t^{\prime}})_{t^{\prime}=1}^{t-L}\in[0,H]^{t-L}\right\}.$
-In other words, the function class $\mathcal{Y}^{t}(\Pi_{S^{t}})$ takes values in the interval $[-1,1]$.
+$\mathcal{Y}^{t}(\Pi_{(S^{t})})\triangleq$$\left\{\frac{y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)}{(L+1)U}:(S^{t^{\prime}})_{t^{\prime}=1}^{t-L}\in[0,H]^{t-L}\right\}.$
+Note that the function class $\mathcal{Y}^{t}(\Pi_{S^{t}})$ takes values in the interval $[-1,1]$.
 We next provide an upper bound on the Pseudoγ-dimension of $\mathcal{Y}^{t}(\Pi_{(S^{t})})$.
@@
 6 Conclusion
-This paper adopts a novel approach by analyzing generalization guarantees of several well-known classes of inventory policies, leveraging the celebrated VC theory.
+This paper analyzes generalization guarantees of several well-known classes of inventory policies on correlated demand sequences, leveraging the celebrated VC theory.
+Our derivation of learning-theoretic dimensions for inventory trajectories diverges from traditional data-driven inventory literature, and our problem-specific analysis for inventory also appears new in the context of data-driven algorithm design.
 Both our theory and simulations reveal important managerial insights into the principle of ?learning less is more?
-in inventory management.
-Depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies.
-Moreover, the number of parameters in a policy class may not accurately measure overfitting.
-For example, the generalization error for $(S^{t})$ policies is comparable to that of $(s,S)$ policies, despite having significantly more parameters.
-Our work has the potential for broader, significant impacts.
-In recent years, many studies have applied state-of-the-art deep learning approaches to solve large-scale inventory problems.
+in inventory management: depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies.
+Moreover, we find surprisingly that $(s,S)$ policies are harder to learn than $(S^{t})$ policies, despite having 2 instead of $T$ parameters.
+Our bound for learning $(S^{t})$ policies based on the powerful VC theory is better than some existing bounds for $(S^{t})$ policies even if they assume independent demands over time.
+All in all, our work advocates for learning well-defined policy classes, which has the potential for broader, significant impacts.
+Indeed, many recent studies have applied state-of-the-art deep learning approaches to solve large-scale inventory problems.
 We believe our structured approach of learning base-stock levels instead of order quantities can potentially enhance existing deep learning techniques, which is particularly valuable in retailing where relevant data is often limited.
@@
 […]
-The authors thank Matias Alvo for providing useful comments on an early draft.
+The authors thank Matias Alvo, Baris Ata, John R.
+Birge, René Caldentey, Xin Chen, Yilun Chen, David A.
+Goldberg, Simai He, Zhiyi Huang, Chengyi Lyu, Hanzhang Qin, Cong Shi, Ellen Vitercik, Amy R.
+Ward, and Sean Zhou for insightful discussions and/or constructive comments on an early draft.